### v2

# MVPatch: More Vivid Patch for Adversarial Camouflaged Attacks on Object Detectors in the Physical World

Abstract
Recent investigations demonstrate that adversarial patches can be utilized to manipulate the result of object detection models. However, the conspicuous patterns on these patches may draw more attention and raise suspicions among humans. Moreover, existing works have primarily focused on enhancing the efficacy of attacks in the physical domain, rather than seeking to optimize their stealth attributes and transferability potential. To address these issues, we introduce a dual-perception-based attack framework that generates an adversarial patch known as the More Vivid Patch (MVPatch). The framework consists of a model-perception degradation method and a human-perception improvement method. To derive the MVPatch, we formulate an iterative process that simultaneously constrains the efficacy of multiple object detectors and refines the visual correlation between the generated adversarial patch and a realistic image. Our method employs a model-perception-based approach that reduces the object confidence scores of several object detectors to boost the transferability of adversarial patches. Further, within the human-perception-based framework, we put forward a lightweight technique for visual similarity measurement that facilitates the development of inconspicuous and natural adversarial patches and eliminates the reliance on additional generative models. Additionally, we introduce the naturalness score and transferability score as metrics for an unbiased assessment of various adversarial patches’ natural appearance and transferability capacity. Extensive experiments demonstrate that the proposed MVPatch algorithm achieves superior attack transferability compared to similar algorithms in both digital and physical domains while also exhibiting a more natural appearance. These findings emphasize the remarkable stealthiness and transferability of the proposed MVPatch attack algorithm.
Index Terms:
Adversarial example, patch attack, physical attack, neural network, transferable and stealthy attack
I Introduction
Figure 1: Introduction to the attack scenarios of MVPatch and illustration of the meaningful and meaningless adversarial patches. (a) demonstrates that MVPatch can make a person invisible to detectors in the real world, while (b) demonstrates that detectors can successfully distinguish a person in the digital world. (c) and (d) demonstrate both diverse meaningful adversarial patches[1] and meaningless adversarial patches[2, 3], respectively.
Deep Neural Networks (DNNs) have achieved remarkable performance in manifold fields, such as computer vision[4, 5, 6, 7], natural language processing[8, 9], and automatic speech recognition[10, 11]. However, due to their lack of interpretability, DNNs are vulnerable to Adversarial Examples (AEs)[12, 13, 14], which raises concerns about their reliability in security-critical applications such as face recognition[15, 16] and autonomous driving[17, 18]. In general, AE attacks can be categorized into two types: Digital Attack (DA)[12, 13, 19, 20, 21, 22, 23, 24, 25, 26], which involves introducing digital perturbations into the input image to carry out the attack, and Physical Attack (PA)[27, 28, 29, 30, 31, 2, 32, 1, 33, 34, 3, 35, 16, 36, 37, 38], which directly targets real-world objects using adversarial perturbations, as illustrated in Fig.1(b) and Fig.1(a), respectively.
Object Detection Models (ODMs)[39, 40, 41], comprised of deep neural layers, are extensively utilized in real-world applications such as people tracking[42], pedestrian and vehicle re-identification (Re-ID)[43, 44], and remote sensing[45]. Although ODMs enhance human convenience and improve the quality of life, they simultaneously pose a considerable risk to individual privacy and sensitive personal information[1, 3, 2, 32]. Furthermore, the capability of physical attacks to be transferred amplifies the vulnerability of ODMs[46, 47, 48]. In real-world scenarios, AEs could be often observed which are quasi-imperceptible to humans. For example, as depicted in Fig.1, an intruder appears under a surveillance camera holding an adversarial patch, which is camouflaged as Van Gogh’s painting, and the DNN-based object detection system will fail. It can even be said that the intruder is invisible under the camera, which will seriously threaten the public safety system Hence, it is imperative to conduct a thorough investigation into PA on ODMs.
Adversarial Patches (APs)[30] have emerged as a potent means to manipulate ODMs within the physical realm, offering several advantages such as input-independence and scene-independence, with substantial real-world impacts[22, 15, 49, 50, 3, 33]. However, much of the extant literature emphasizes enhancing the efficacy of the attack within this physical sphere, rather than seeking to optimize its stealth attributes and transferability potential. Hu et al.[1] considered how to improve the stealthiness of the adversarial patch by Generative Adversarial Network (GAN)[51]. In contrast, Guesmi et al.[38] proposed a lightweight approach to generate natural patches without GAN. Moreover, many scholars have endeavored to augment APs’ transferability. A case in point is the research by Wang et al.[46], who honed AP transferability within automated checkout systems through a bias-guided arrangement. In a similar endeavor, Wei et al.[47] devised a Pay No Attention (PNA) attack, wherein attention gradients are disregarded during backpropagation, thereby effectively improving backpropagation’s linearity. Furthermore, Huang et al.[48] modified a single surrogate model into a diverse ensemble structure, yielding more transferable adversarial examples. Nevertheless, numerous challenges persist, which include the substantial computational resource requirements associated with the use of GAN during the training process, and an unrelenting research focus on enhancing the stealthiness[2, 31, 52, 1, 32, 3, 35, 29, 53, 28, 38] and transferability[46, 47, 48] of APs. Significantly, there remains limited research simultaneously addressing both stealthiness and transferability issues related to APs, with an additional focus on computational resource efficiency.
To bridge this gap, we propose a dual-perception-based framework to generate a potent adversarial patch, referred to as the MVPatch (More Vivid Patch). This attack framework offers remarkable stealthiness and transferability capacities for attacking against ODMs. To improve the transferability potential, we present a model-perception degradation method for ensemble multiple object detectors achieved through the ensemble attack loss function. Additionally, to enhance the stealthiness attributes, we present a human-perception improvement method realized by the Compared Specified Image Similarity (CSS) loss function, which allows for the generation of natural and stealthy adversarial patches without the reliance on additional generative models. By integrating both the model-perception-based method and the human-perception-based method, we can generate a more powerful adversarial patch with significant transferable attack ability and invisibility. Moreover, we introduce the naturalness score and the transferability score as experimental evaluation metrics to assess the attack performance regarding stealthiness and transferability. We exploit the YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny as threat models and Faster RCNN, SSD, and YOLOv5 as transfer attack models. Based on extensive experiments conducted in both digital and physical scenarios, along with several independent subjective surveys, our proposed MVPatch consistently generates more natural and transferable results compared to several state-of-the-art baselines. We present the primary specialization of MVPatch, along with the distinguishing features of adversarial patches that possess meaning and those that are devoid of meaning. The main contributions of this work can be summarized as follows:
  • To the best of our knowledge, we are the first to propose a dual-perception-based attack framework for generating an adversarial patch (MVPatch). This patch offers both stealthiness and transferability by exploiting both model-perception-based and human-perception-based methods.
  • We present the model-perception-based method, an attack method that reduces the object confidence scores of an ensemble of multiple object detectors using an ensemble attack loss function. This enhances the transferability of the adversarial patches, significantly improving attack performance in both digital and physical domains compared to state-of-the-art baselines.
  • We propose a human-perception-based method, a lightweight visual similarity measurement method that generates natural and stealthy adversarial patches without the need for additional generative models. This allows us to generate patches based on specific images.
  • We employ naturalness and transferability scores as evaluation metrics to assess the naturalness and attack transferability of the diverse adversarial patches in our experiment.
  • We perform a comprehensive analysis of the transfer attack performance and naturalness of the proposed method, using both meaningful and meaningless patch approaches in a variety of digital and physical scenarios.
The remainder of this paper is organized as follows. Section II provides a concise literature review on adversarial examples and adversarial patches in various scenarios, encompassing both digital and physical contexts. In Section III, we introduce the definition and characteristics of adversarial patch attacks, along with detailed information regarding the input and output of the threat models. The generation process of our proposed MVPatch is then presented in this section. Section IV introduces the environment and dataset in the experiment, followed by the provision of relevant experimental details. Additionally, we propose evaluation metrics for the experiment. Section V presents empirical evidence that showcases the effectiveness of the proposed MVPatch through comprehensive evaluations conducted in both digital and physical environments. Ultimately, we summarize the paper and present the further research directions in Section VI.
II Related Work
In this section, we briefly review the literature related to adversarial examples and adversarial patches in diverse scenarios, such as in the digital domains and physical domains. Furthermore, we discuss the recent issues to be solved and present the necessity of our work.
II-A Digital Attack
Adversarial Examples (AEs) are elaborately designed perturbations that are imperceptible to humans but could mislead Deep Neural Networks (DNNs). In 2014, Szegedy et al.[12] discovered a special property of DNNs known as Adversarial Examples (AEs). They successfully attacked a DNN using the L-BFGS algorithm, causing it to produce incorrect predictions. In 2015, Goodfellow et al.[13] introduced the Fast Gradient Sign Method (FGSM), which improved the effectiveness and success rates of attacks on neural network classification models. Kurakin et al.[22] developed the Basic Iterative Method (BIM) and demonstrated the application of AEs in the physical world. Papernot et al.[23] focused on partial attack and presented the Jocobian-Based Saliency Map Attack (JSMA) algorithm. Madry et al.[19] proposed Project Gradient Descent (PGD), an innovative framework that combines attack and defense algorithms to enhance the robustness of DNNs. There are numerous varieties of adversarial attacks in the digital domain, e.g., Deepfool[21], C&W[20], ZOO[24], Universal Perturbation[25], One-Pixel attack[26] and some attacks on large language vision models[54, 55, 56].
II-B Physical Attack
Recently, it has been observed that printed AEs can effectively deceive neural network models in the physical domain. Athalya et al.[27] developed robust three-dimensional AEs that can mitigate the impact of diverse angle transformations. Sharif et al.[28] successfully attacked facial recognition systems by constructing adversarial eyeglasses, while Komkov et al.[29] proposed the construction of an adversarial hat for the same purpose. In 2018, Brown et al.[30] introduced the concept of Adversarial Patch (AP). Liu et al.[31] proposed DPatch to deceive object detection models, while at around the same time, Thy et al.[2] introduced AdvPatch to mislead automated surveillance systems. Wu et al. [32] presented the AdvCloack to make humans invisible in the object detector. Hu et al.[1] designed a more natural patch based on the AdvPatch[2]. Xu et al.[33] proposed to construct an adversarial T-shirt to mislead the object detection systems. Huang et al.[34] introduced the Transfer-based Self-ensemble Attack (T-SEA) for attacking object detectors using ensemble attacks, effectively improving the transferable attack performance. Hu et al.[3] proposed the Toroidal-Cropping-based Expandable Generative Attack (TC-EGA) algorithm, which applies adversarial textures to clothing to improve the stealthiness of the adversarial patch. Wang et al.[35] constructed the adversarial patch based on the attention mechanism for the car camouflage. Wei et al.[16] investigated methods to enhance the attack performance of adversarial patches by focusing on their spatial positioning. Liu et al.[36] proposed a Perceptual-Sensitive Generative Adversarial Network (PS-GAN) to enhance the visual fidelity and attacking performance for the adversarial patches. Xue et al.[37] presented the Diffusion-Based Projected Gradient Desent (DIff-PGD) to generate realistic adversarial examples by using diffusion models. Guesmi et al.[38] proposed AdvART to attack object detectors by a simple method without generative models.
Despite the significant results achieved by prior research on adversarial patches in the physical world, there are still some issues to be addressed, such as the trade-off between transferability[34, 1, 57] and stealthiniess[27, 29, 32, 33, 3, 35, 58, 36, 37, 38] of the adversarial patch. Furthermore, most recent research focuses on generating meaningless adversarial patches. The meaningless adversarial patch means that has no sense for human eye perception as Fig.1(d) illustrates. To the best of our knowledge, only one study has systematically examined the investigation of meaningful and transferable adversarial patches through the utilization of a generative adversarial network (GAN) by Hu et al.[1]. However, training a generative adversarial network is challenging and computationally expensive. To address these issues, we propose the MVPatch algorithm, which outperforms other state-of-the-art baseline methods in terms of transferable and stealthy attacks both in the digital and physical scenarios while maintaining a lightweight approach.
III Methodology
Figure 2: The MVPatch pipeline involves embedding masks into benign images and applying them to object detectors to determine the object confidence scores. To achieve improved attack performance, the algorithm minimizes various losses, such as TV loss, NPS loss, OBJ loss, and CSS loss, to obtain the optimal adversarial patch. Algorithm 1 illustrates the complete procedure for the MVPatch algorithm.
In this section, we first introduce the definition and characteristics of adversarial patch attacks and then detail the input and output information of threat models. Finally, we summarize the entire generation process of MVPatch in this section.
III-A Problem Formulation
The objective of this study is to generate a malicious image $\mathcal{I^{*}}$ with an adversarial patch from a benign image $\mathcal{I}$, to deceive object detection models. The complete process of generating the adversarial patch is illustrated in Fig.2. Typically, the malicious images with adversarial patches can be formulated as:
*$\mathcal{I^{*}}=(1-\mathcal{M})\odot\mathcal{I}+\mathcal{M}\odot\mathcal{P},$ (1)
where $\mathcal{M}$ represents the patch mask, which determines the size, shape, and location of the adversarial patch $\mathcal{P}$. The $\odot$ symbol denotes the Hadamard product, which multiplies the elements of the corresponding positions in the matrices $\mathcal{M}$, $\mathcal{I}$, and $\mathcal{P}$. We denote the object detection model as $f(x):x\in\mathbb{R}^{\mathbf{H\times W\times C}}\mapsto\mathcal{Y}\in\mathbb{R}^{6}$. $x$ is input image and $\mathcal{Y}$ is label of object. The generation of adversarial patches can be viewed as a constrained optimization problem:
$f(\mathcal{I^{*}})\neq f(\mathcal{I})\quad\mathrm{s.t.}\,\lVert\mathcal{I^{*}}-\mathcal{I}\rVert\longrightarrow\mathcal{P},$ (2)
where $\mathcal{P}$ represents the adversarial patch generated under the constraints of *$\lVert\mathcal{I^{*}}-\mathcal{I}\rVert\longrightarrow\mathcal{P}$, while satisfying $f(\mathcal{I^{*}})\neq f(\mathcal{I})$. Finding a closed-form solution for this optimization problem is not feasible due to the non-convex nature of the deep neural network model $f(\cdot)$. Therefore, to approximately solve this optimization problem, we utilize the following equation:
$\underset{\mathcal{P}}{\mathrm{argmax}}\,\mathcal{L}(f(\mathcal{I^{*}}),f(\mathcal{I})),$ (3)
where $\mathcal{L}$ represents the cross-entropy loss function. To solve the optimization problem, we can utilize the Adam optimizer, which is one of the most common and widely used optimization techniques.
III-B Threat Models
We apply the YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny models to generate adversarial patches. These patches are then evaluated on the YOLOv5, Faster-RCNN, and SSD models to assess their transferability in attacks. In the object context of the YOLO series, given the benign image $\mathcal{I}$, the objective of the adversarial attack is to jeopardize the object detector $f_{\theta}(x,\mathcal{Y})$, where $x$ belongs to $\mathcal{I}$, $\mathcal{Y}=\{\mathrm{a},\mathrm{b},\mathrm{w},\mathrm{h},\mathrm{obj},\mathrm{cls}\}$ and $\theta$ denotes the parameter of the object detector. We suppress the object confidence score by utilizing the adversarial patches. $\mathrm{a}$ and $\mathrm{b}$ are represented by the value of the horizontal and vertical coordinates of the detection frame. $\mathrm{w}$ and $\mathrm{h}$ are the width and height of the detection frame. $\mathrm{obj}$ and $\mathrm{cls}$ are object confidence scores and classification scores. Our focus is to suppress the person’s object confidence score by optimizing the use of adversarial patches.
III-C Robustness of Adversarial Patch in the Physical Domain
Previous studies have primarily focused on applying adversarial patches in the digital domain. However, it is important to recognize that adversarial patches can also appear in the physical domain. To enhance the effectiveness of adversarial patches in physical attacks, we propose the utilization of the Non-Printable Score (NPS)[28] and Total-Variation (TV)[3] as loss functions. These will aid in the creation of more aggressive and robust adversarial patches specifically designed for physical scenarios. We present the NPS for improving the robustness when we generate adversarial patches in the physical domain using digital patterns and to ensure compatibility with the iPad color gamut and printer gamut. The formulation of the NPS is as follows:
$\mathcal{L_{\mathrm{NPS}}}=\sum_{c\in\mathcal{C},p\in\mathcal{P}}|p-c|,$ (4)
where $\mathcal{P}$ and $\mathcal{C}$ represent the pixel values of the adversarial patch and the colors that the iPad can display, respectively. Both $p$ and $c$ are elements of $\mathcal{P}$ and $\mathcal{C}$. To improve the physical robustness of adversarial patches, we utilize the Euclidean distance ||$\cdot$|| ⋅ | to constrain the generation of adversarial patches.
We also introduce the TV term, which enhances the smoothness of neighboring pixels. The TV formulation is defined as:
$\mathcal{L_{\mathrm{TV}}}=\sum_{a,b}|p_{a,b}-p_{a,b+1}|+|p_{a,b}-p_{a+1,b}|,$ (5)
where $p$ denotes the pixel values of the image, while $a$ and $b$ represent the abscissa and ordinate values of a single pixel, respectively.
III-D Human-perception-based Method
There is a trade-off between the naturalness and attack performance of adversarial patches. For instance, if an adversarial patch exhibits higher attack performance, it tends to have lower naturalness. In other words, as the similarity to the source image decreases, the success rate of the attack increases. To obtain a more natural adversarial patch, we propose the human-perception-based method by utilizing the Minkowski distance as a loss function to constrain the naturalness of the patch, which we refer to as the Compared Specified Image Similarity (CSS) measurement. The CSS is formulated as follows:
$\mathcal{L_{\mathrm{CSS}}}=\left(\sum_{i=1}^{N}|\hat{x_{i}}-x_{i}|^{p}\right)^{\frac{1}{p}},$ (6)
where $\hat{x}$ represents the adversarial patch generated using the specified image $x$. We choose $p=2$ in this work and $i$ represents the number of pixels of $\hat{x}_{i}$ and $x_{i}$.
III-E Model-perception-based Method
To improve the transferability of adversarial patches, we propose the model-perception-based method. First, we present the use of two distinct loss functions to attack object detection models, as shown below:
$\mathcal{L}_{\mathrm{OBJ_{1}}}=\frac{1}{N}\sum_{i=1}^{N}S^{i}_{\mathrm{OBJ}}(\mathcal{I^{*}},\theta),$ (7)
$\mathcal{L}_{\mathrm{OBJ_{2}}}=\mathop{\max}_{\mathcal{I^{*}}}(\{S^{i}_{\mathrm{OBJ}}(\mathcal{I^{*}},\theta)\}^{N}_{i=1}),$ (8)
where $S^{i}_{\mathrm{OBJ}}$ denotes the object confidence score of the $i^{th}$ object detection model in the ensemble of object detection models (YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny). $\mathcal{I^{*}}$ is the image with adversarial patch and $\theta$ corresponds to the parameters of the object detection model.
TABLE I: Attack Performance and Comprehensive Evaluation with Diverse Objective Loss Functions
Faster-RCNN SSD YOLOv2 YOLOv3 YOLOv3-tiny YOLOv4 YOLOv4-tiny YOLOv5 Mean mAP CS ED
$\mathcal{L}_{\text{OBJ}_{1}}$ 42.07% 54.57% 30.26% 31.39% 13.49% 27.31% 22.74% 35.72% 32.20% 97.91% 0.0428
$\mathcal{L}_{\text{OBJ}_{2}}$ 49.40% 61.08% 50.34% 37.47% 41.44% 52.64% 56.46% 56.46% 51.80% 97.91% 0.0419
We compare the attack performance of diverse object confidence score loss functions on various object detection models. It is observed that the adversarial patch generated by $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ outperforms the one generated by $\mathcal{L}_{\mathrm{{OBJ}_{2}}}$ in terms of attacking the object detection model. The adversarial patches generated through diverse loss functions are then applied to eight object detection models (Faster-RCNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5 models) to assess their attack performance and similarity to specified images. The adversarial patch generated by $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ not only exhibits excellent attack capability (the average mAP of $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ is 32.20% and the average mAP of $\mathcal{L}_{\mathrm{{OBJ}_{2}}}$ is 51.80%) but also maintains a high level of naturalness. Consequently, we select $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ as our objective loss function. The results of the experiment are illustrated in Table I, where the Faster-RCNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5 models are included as threat models. The mean mAP represents the average value of the mean Average Precision (mAP) obtained from diverse models. CS corresponds to the Cosine Similarity and ED denotes the Euclidean Distance.
III-F Dual-perception-based Framework to Generate MVPatch
To make the optimization process converge faster, we introduce additional optimization factors to the previously mentioned loss functions and formulate an ensemble loss function as Equation (9):
$\mathcal{L_{\mathrm{TOTAL}}}=\alpha\mathcal{L_{\mathrm{NPS}}}+\beta\mathcal{L_{\mathrm{TV}}}+\lambda\mathcal{L_{\mathrm{CSS}}}+\mathcal{L_{\mathrm{OBJ}}}.$ (9)
Since optimizing the ensemble loss function with the Adam optimizer is challenging, we modified our optimization strategy. Instead of reducing the learning rate as training epochs increase, we set thresholds for the parameters and decrease the learning rate when a parameter no longer exhibits optimization during the training process. The updating strategy for the learning rate is governed by Equation (10):
$\mathcal{LR_{\mathrm{N}}}=\gamma\mathcal{LR_{\mathrm{O}}},$ (10)
where $\mathcal{LR}_{\mathrm{N}}$ is the new learning rate, while $\mathcal{LR}_{\mathrm{O}}$ corresponds to the old learning rate. In our experiments, we utilize $\gamma=0.01$ as the decay coefficient.
The complete process is summarized in Algorithm 1. This algorithm aims to generate an adversarial patch, referred to as MVPatch, for a given original image $\mathcal{I}$ using a neural network $f(x)$ with the output being the adversarial patch $\mathcal{P}$.
Algorithm 1 Generating Adversarial Patch (MVPatch)
1:  Input: Original Image $\mathcal{I}$, Neural Network $f(x)$
2:  Output: Adversarial Patch $\mathcal{P}$
3:  Initialize patch mask $\mathcal{M}$
4:  Initialize patch $\mathcal{P}$ using $\mathcal{M}$
5:  *$\mathcal{I^{*}}=(1-\mathcal{M})\odot\mathcal{I}+\mathcal{M}\odot\mathcal{P}$
6:  $\mathcal{L}_{\mathrm{\text{TOTAL}}}\leftarrow\alpha\mathcal{L}_{\mathrm{\text{NPS}}}+\beta\mathcal{L}_{\mathrm{\text{TV}}}+\lambda\mathcal{L}_{\mathrm{\text{CSS}}}+\mathcal{L}_{\mathrm{\text{OBJ}}}$
7:  Initialize learning rate $\mathcal{LR}$ and epochs $E$
8:  while not converged or $E<E_{\mathrm{\text{MAX}}}$ do
9:     Update $\mathcal{I^{*}}$ using backpropagation based on $\mathcal{L}_{\mathrm{\text{TOTAL}}}$
10:     Update patch mask $\mathcal{M}$ based on $\mathcal{I^{*}}$
11:     Update patch $\mathcal{P}$ based on $\mathcal{M}$
12:     $E\leftarrow E+1$
13:     if parameter optimization threshold not reached then
14:        $\mathcal{LR}\leftarrow\gamma\mathcal{}$
15:     end if
16:  end while
17:  Return: Adversarial Patch $P$
IV Experiments
In this section, we briefly introduce the environment and dataset in the experiment. Additionally, we propose evaluation metrics for the experiment. In the digital environment, we generate meaningful and meaningless adversarial patches and compare them with similar patch attack approaches. In the physical environment, we compare the Attack Success Rate (ASR) of MVPatch with similar patch attack methods and vary the angles and distances between the patch and camera to investigate their impacts on the ASR.
IV-A Implementation Details
In the digital setting, we initialize the adversarial patch using Gaussian noise. We then segment the patch mask onto the person image and overlay the adversarial patch onto the mask. Finally, we input the image with the adversarial patch into the object detection model. The threat models considered in our experiments are YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny. For evaluating the transferability of the attack, we utilize YOLOv5, Faster-RCNN, and SSD as the transferable attack models. Subsequently, we calculate the loss functions, including $\mathcal{L}_{\mathrm{OBJ}}$, $\mathcal{L}_{\mathrm{NPS}}$, $\mathcal{L}_{\mathrm{CSS}}$, and $\mathcal{L}_{\mathrm{TV}}$. To optimize the overall loss function $\mathcal{L}_{\mathrm{TOTAL}}=\alpha\mathcal{L}_{\mathrm{NPS}}+\beta\mathcal{L}_{\mathrm{TV}}+\lambda\mathcal{L}_{\mathrm{CSS}}+\mathcal{L}_{\mathrm{OBJ}}$ using backpropagation and the Adam optimizer, we set parameters as $\alpha=0.01$, $\beta=2.5$, and $\lambda=2.5$.
In the physical setting, we employ YOLOv2 as the experimental model to evaluate the Attack Success Rate (ASR) of the generated patches. The parameters of YOLOv2, including $\text{Conf}_{\text{THRESHOLD}}=0.75$ and $\text{NMS}_{\text{THRESHOLD}}=0.5$, are carefully selected. For collecting the source images, we utilize the iPhone 11 camera, which serves as our image input device.
IV-B Experimental Environment and Dataset
For training and evaluating the proposed approach, we utilize the Inria Person dataset[59], which is specifically designed for pedestrian detection tasks. The dataset comprises 614 images for training and 288 images for testing. We utilize an RTX 3090 for the computational module and a $12^{*}$ Xeon Platinum 8260C for task scheduling. Python version 3.6 and PyTorch version 1.6.1 are used.
In the digital setting, we evaluate the performance of the proposed approach using meaningful adversarial patches and meaningless adversarial patches. These patches are added to images of people from the Inria dataset.
In the physical setting, we print the generated MVPatch using an iPad and attach it to the chest of the investigator. By varying the angles ($0^{\circ}$, $30^{\circ}$, $60^{\circ}$, $90^{\circ}$) and distances (1 meter, 2 meters, 3 meters, 4 meters) between the human and the camera, we investigate the attack performance of MVPatch and other similar attack methods in a physical setting.
IV-C Evaluation Metrics
To evaluate the naturalness and attack transferability of MVPatch, we employed the following evaluation metrics: mean Average Precision (mAP), Naturalness Score (NS), and Transferability Score (TS).
mAP (mean Average Precision): The mAP is a widely used metric calculated by summing the average precision of models for each test sample and dividing it by the total number of test samples. It allows for comparisons of model performance across different datasets. The precision PREC is computed as the ratio of true positive samples divided by the sum of true positive and false positive samples. The recall REC is calculated as the ratio of true positive samples divided by the sum of true positive and false negative samples. Formally,
$\begin{split}\mathrm{PREC}&=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}},\quad\mathrm{REC}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}},\\ \mathrm{mAP}&=\frac{1}{n-1}\sum_{i=1}^{n-1}(\mathrm{REC}_{i+1}-\mathrm{REC}_{i})\mathrm{PREC}_{i+1},\end{split}$ (11)
where TP, FP, TN, and FN denote respectively true positive sample, false positive sample, true negative sample, and false negative sample. Through the equation mentioned above, we can obtain the mAP of the object detection models.
NS (Naturalness Score): The NS is used to measure the similarity between the adversarial patch and the specified image, which is defined as
$\mathrm{NS}=$ $\lambda\left(\frac{\mathrm{CSS}-\widehat{\mathrm{CSS}}_{\mathrm{NG}}}{\widehat{\mathrm{CSS}}_{\mathrm{S}}-\widehat{\mathrm{CSS}}_{\mathrm{NG}}}+\frac{\mathrm{CSS}-\widehat{\mathrm{CSS}}_{\mathrm{NR}}}{\widehat{\mathrm{CSS}}_{\mathrm{S}}-\widehat{\mathrm{CSS}}_{\mathrm{NR}}}\right)$ (12)
$+\left(1-\lambda\right)\left(\frac{\mathrm{ED}-\widehat{\mathrm{ED}}_{\mathrm{NG}}}{\widehat{\mathrm{ED}}_{\mathrm{NG}}-\widehat{\mathrm{ED}}_{\mathrm{S}}}+\frac{\mathrm{ED}-\widehat{\mathrm{ED}}_{\mathrm{NR}}}{\widehat{\mathrm{ED}}_{\mathrm{NR}}-\widehat{\mathrm{ED}}_{\mathrm{S}}}\right),$
where CSS represents the cosine similarity score between the specified image and adversarial patch and $\widehat{\text{CSS}}$ is the cosine similarity score between the specified image and adversarial patch of NG and NR images. ED denotes the Euclidean distance. NG and NR are gray-scale and random noise. S is an abbreviation for the source image. A higher naturalness score indicates that MVPatch has a stronger similarity to the specified image.
TS (Transferability Score): The TS measures the transferable attack performance of the adversarial patch across diverse object detection models, which is defined as
$\mathrm{TS}=\frac{1}{N}\sum_{i=0}^{N}\left(\frac{\|D_{i}-\widehat{D}_{i}^{\mathrm{NG}}\|_{2}}{\widehat{D}_{i}^{\mathrm{NG}}}+\frac{\|D_{i}-\widehat{D}_{i}^{\mathrm{NR}}\|_{2}}{\widehat{D}_{i}^{\mathrm{NR}}}\right),$ (13)
where $D_{i}$ denotes the mAP of object detection models detecting normal images while $\widehat{D}_{i}$ represents the mAP of object detection models detecting NG and NR images. NG and NR represent gray-scale and random noise. $||\cdot||_{2}$ is the L2 norm to constrain the distances between $D_{i}$ and $\widehat{D}_{i}$. The higher transferability score reveals that the adversarial patch has a higher transferable attack performance. We define
$\mathrm{ASR}=\frac{1}{|\mathcal{D}_{\mathrm{test}}|}\sum_{\mathcal{I}\in\mathcal{D}_{\mathrm{test}}}\mathbbm{1}\{f(\mathcal{I^{*},\theta})\neq f(\mathcal{I},\theta)\},$ (14)
as an indicator of the attack success rate, where $f(\cdot)$ represents the consequence of the models’ detection. $\mathcal{D}_{\mathrm{test}}$ denotes a set of benign images from test datasets. The ASR equals 1 when $f(\mathcal{I^{*},\theta})\neq f(\mathcal{I},\theta)$, and equals 0 otherwise.
V Results
This section presents empirical evidence that showcases the effectiveness of the proposed MVPatch through comprehensive evaluations conducted in both digital and physical domains. In the digital domain, we divided the experimental objects into meaningful adversarial patches and meaningless adversarial patches. In the physical domain, we assessed the patch attack performance from various angles and distances by utilizing a camera to capture images as input sources. The experimental results demonstrate that our method outperforms other similar adversarial patch attack methods, both in the digital and physical domains. In the last, we show the visual consequence of the experiment in the digital domain and physical domain.
V-A Evaluation of Naturalness in the Digital World
To assess the naturalness of the adversarial patch, we utilize the YOLO series models as threat models. The natural factor, represented by the parameter $\lambda$, is adjusted within the range of [1,3] to generate a variety of adversarial patches. The naturalness score (NS) is then calculated to quantify the naturalness of these patches. The generated adversarial patches with different naturalness scores are displayed in Fig.3. As we move from left to right on the coordinate axis, the naturalness score gradually increases, indicating a higher level of naturalness.
TABLE II: Experimental Consequence of Naturalness of MVPatch
Adversarial Patches Evaluation Metrics
mAP CS ED NS
SRC 65.70% 100% 0.00 100.00
NoiseGrey 74.05% 97.45% 6.01 0.00
NoiseRandom 75.09% 92.82% 8.27 0.00
$\lambda=1$ 32.30% 97.61% 4.63 35.59
$\lambda=1.25$ 32.19% 97.87% 4.24 42.11
$\lambda=1.5$ 34.39% 98.54% 3.73 56.77
$\lambda=1.75$ 33.13% 98.78% 3.44 62.48
$\lambda=2$ 33.98% 99.01% 3.12 68.14
$\lambda=2.25$ 37.49% 99.20% 2.82 72.97
$\bm{\lambda=2.5}$ 36.63% 99.30% 2.53 76.08
$\lambda=2.75$ 37.46% 99.38% 2.21 78.94
$\lambda=3$ 39.69% 99.44% 1.99 81.01
Explanation for Experimental Consequence: $\lambda=[1,3]$ represents the adversarial patches generated by diverse natural factors. SRC represents the source image. NoiseGrey and NoiseRandom are matrices with values of 0.5 and random values, respectively. We deploy the mAP (mean Average Precision), CS (Cosine Similarity), and ED (Euclidean Distance) to measure the attack performance and the similarity of the adversarial patch with the source image.
Figure 3: The naturalness score (NS) of adversarial patches with diverse $\lambda$ parameters. As the NS increases, the level of similarity between the source image and the generated image rises.
Analysis and Conclusion for Experimental Consequence: As shown in Table II, we vary $\lambda$ from 1 to 3, resulting in an increase in the score of naturalness from 35.59 to 81.01. Simultaneously, the mAP increases from 32.30% to 39.69%. However, a higher mAP generally corresponds to greater accuracy of the object detection models and a lower attack performance of adversarial patches. There is a tradeoff between naturalness and attack performance, so it is not possible to obtain an adversarial patch with both higher naturalness and attack performance. We identify a threshold for the score of naturalness that helps us obtain the optimal adversarial patch with both naturalness and attack performance. Through an analysis of 20 experimental participants, it is generally agreed that when the score of naturalness is above 75 ($\lambda=2.5$), the generated adversarial patch would be considered more natural. Furthermore, if we continue to increase the natural factor $\lambda$, the gain in naturalness recognizable to the human eye will decrease. Therefore, we define $\lambda=2.5$ as the parameter of the loss function in the following experiments. The images with different scores of naturalness for the adversarial patches are shown in Fig.3.
TABLE III: Experimental Consequence of Transferability of Meaningful Adversarial Patches (mAP %)
${}^{(MeaningfulAdversarialPatches)}$hreat Models Transfer Attack Models Black Box(TS $\uparrow$)
Faster RCNN SSD YOLOv2 YOLOv3 YOLOv3-tiny YOLOv4 YOLOv4-tiny YOLOv5
[…]
$\bm{{}^{\mathrm{(-)}}}$nsemble Model[1] 61.28 52.28 49.42 35.46 25.29 51.71 18.51 64.00 40
[…]
Figure 4: Adversarial patches are generated by various object detectors, including YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny and compared attack performance on the individual model with Natural Patch[1]. Additionally, we compare the performance of our ensemble attack with that of the Natural Patch. The attack performance of diverse adversarial patches are illustrated as Table III.
TABLE IV: Experimental Consequence of Transferability of Meaningless Adversarial Patches (mAP %)
[…]
Faster-RCNN SSD YOLOv2 YOLOv3 YOLOv3-tiny YOLOv4 YOLOv4-tiny YOLOv5
AdvPatch[2] 43.11 48.57 4.69 47.59 34.24 58.74 33.82 49.65 46.29
AdvTexture[3] 48.18 36.12 5.31 46.83 18.74 63.03 34.13 60.47 47.56
AdvCloak[32] 55.19 60.82 33.74 54.77 53.42 67.57 56.12 68.05 24.62
MVPatch[Ours] 34.95 48.95 23.93 13.09 6.04 26.72 14.77 34.35 64.47
TABLE V: Experimental Consequence of MVPatch and other Patches in the Physical Domain
Adversarial Patches Evaluation Metrics
ASR Number of Images Stealthiness Attack Physical Attack Transferable Attack Lightweight Model
[…]
Figure 5: Adversarial patches, such as AdvPatch[2], AdvTexture[3], Natural Patch[1], and our MVPatch, are employed in the physical world using diverse attack methods. TableV displays the attack success rate of different adversarial patches, as well as other important evaluation metrics.
(a) Impact of diverse angles to ASR of adversarial patches
(b) Impact of diverse distances to ASR of adversarial patches
Figure 6: Experimental consequence illustration of the adversarial patches attacking in the physical domain. Fig.6(a) illustrates the impact of varying camera angles on the ASR of different attack methods, such as Natural Patch, AdvPatch, AdvTexture, Benign Image, and MVPatch. Fig.6(b) illustrates the impact of varying distances between the camera and patches on the ASR of different attack methods, such as Natural Patch, AdvPatch, AdvTexture, Benign Image, and MVPatch.
V-B Evaluation of Attack Transferability in the Digital domain
To systematically investigate the attack performance of adversarial patches in the digital domain, we divide the experimental objects into two parts: meaningful adversarial patches and meaningless adversarial patches. Most research mainly focused on meaningless adversarial patches and few researchers investigated the meaningful and meaningless adversarial patches in the digital and physical domains.
V-B1 Evaluation of Attack Transferability of Meaningful Adversarial Patches in the Digital domain
We utilize the pre-trained models including YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny to generate diverse adversarial patches with meaningful content using the natural factor $\lambda=2.5$, as shown in Fig.4, 4, 4, 4, and 4. By combining these pre-trained models, we obtain ensemble attack adversarial patches, as shown in Fig.4 and 4. To compare the influence of person images and other images as adversarial patches for the object detection models, we use the sunflower image (Fig.4) as a control test for the person image (Fig.4). We input all the acquired adversarial patches into transfer attack models, including Faster R-CNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5, to calculate the transferability score and compare it with the Natural Patch[1]. The comparable experimental results are shown in Table III.
Explanation for Experimental Consequence: YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, YOLOv5, Faster R-CNN, and SSD are pre-trained object detection models. The ”Threat Models” column represents diverse adversarial patches generated by different models, including YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and the Ensemble Model, which combines YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny. The ”original image” refers to an image without an adversarial attack.”Grey-scale Noise” and ”Random Noise” separately indicate an image with padding of a value of 0.5 and an image with random values. The Black Box (TS) measures the transferability of adversarial attacks, with the transferability score (TS) indicating superior performance with a higher score.
Evaluation and Analysis for Experimental Consequence: Firstly, we compare the transferability score of MVPatch (Ours) to that of Natural Patch[1] on individual object detection models. MVPatch performs better than Natural Patch on most of the models in the experiment. For the YOLOv4 model, the TS of MVPatch is approximately 33% higher than the TS of Natural Patch. Similarly, the excellent attack performance of MVPatch is observed on the other four object detection models (YOLOv2, YOLOv3, YOLOv3-tiny, and YOLOv4-tiny). On the other hand, the attack performance of MVPatch on the ensemble model is surpassed by Natural Patch, with the TS of MVPatch on the ensemble model being approximately 16% higher than the TS of Natural Patch.
From the results of the meaningful adversarial patch experiment, it can be seen that the proposed algorithm for meaningful adversarial patches, MVPatch, achieves superior performance compared to the similar algorithm, Natural Patch, in terms of attack transferability.
V-B2 Evaluation of Attack Transferability of Meaningless Adversarial Patches in the Digital domain
We adjust the natural factor of $\lambda=0$ and utilize the joint pre-trained models, including YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny, to generate meaningless adversarial patch MVPatch. In comparison, we consider AdvPatch[2], AdvTexture[3], and AdvCloak[32] as subjects for the experiment on meaningless adversarial patches. We use all the obtained adversarial patches as input for Faster R-CNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5, calculate the transferability scores, and compare them with those of AdvPatch [2], AdvTexture[3], and AdvCloak[32]. The comparable experimental results are presented in TableIV.
Explanation for Experimental Consequence: YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, YOLOv5, Faster R-CNN, and SSD are pre-trained object detection models. The ”Meaningless Adversarial Patches” column represents various meaningless adversarial patches generated by diverse attack algorithms. The Black Box (TS) measures the transferability of adversarial attacks, with a higher transferability score indicating better performance. TS stands for the transferability score.
Evaluation and Analysis for Experimental Consequence: As illustrated in Table IV, the transferability score (TS) of MVPatch is approximately 40% higher than that of AdvCloak. Compared to AdvPatch and AdvTexture, the TS of MVPatch is approximately 20% higher. Moreover, when the transfer models, including YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, YOLOv5, Faster R-CNN, and SSD, are attacked by MVPatch, their mean Average Precision (mAP) reduces by 10 to 20%, compared to the other three algorithms.
To sum up, in terms of meaningless adversarial patches, the proposed algorithm, MVPatch, exhibits higher attack transferability compared to similar algorithms.
V-C Evaluation of Attack Performance of Adversarial Patches in the Physical domain
To systematically investigate the attack performance of adversarial patches in the physical domain, we divide the experimental conditions into two parts: diverse angles and diverse distances. Prior research mainly focused on the attack performance of adversarial patches in the physical domain and few researchers investigated how impact the attack performance of adversarial patches using diverse angles and distances.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(a) Benign Images
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(b) Malicious Images
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(c) Benign Images
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(d) Malicious Images
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(e) Benign Images
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(f) Malicious Image
Figure 7: Illustration of experimental consequence in the digital domain. The images with green rectangles represent the object detectors can recognize the class of a person successfully while the images with red rectangles denote the object detectors can not recognize the class of a person.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(a) MVPatch
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(b) MVPatch
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(c) Natural Patch
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(d) Natural Patch
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(e) Adv. Patch
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(f) Adv. Texture
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(g) Benign Image
Figure 8: This illustration demonstrates the experimental outcomes in the physical domain. Images with a green rectangle represent successful recognition of the person’s class by the object detectors, whereas images with a red rectangle indicate the object detectors’ failure to recognize the person’s class. The images at the bottom are adversarial patches and mosaics are added after detection, to avoid privacy disclosure for volunteers.
V-C1 Evaluation of Attack Performance of Diverse Adversarial Patches in the Physical Domain
The volunteers hold various adversarial patches, including AdvPatch, AdvTexture, a benign image, Natural Patch, and the proposed MVPatch, against their chests. The patches are positioned at an angle of {$0^{\circ},30^{\circ},60^{\circ},90^{\circ}$} and a distance of {1, 2, 3, 4} meters. We use an iPhone 11 to capture the images with the adversarial patches from diverse distances and angles in the physical world. In total, we collected 9681 images. To preserve the pixel details of the adversarial patches, we view them using an iPad Pro. The attack success rates and some evaluation metrics of different adversarial patches are shown in Table V, and sample patches are displayed in Fig.5.
Explanation for Experimental Consequence:The AdvPatch[2], and AdvTetxure[3] are both meaningless adversarial patches and the Natural Patches[1] and proposed MVPatches are meaningful patches. The Benign Image refers to an image without adversarial examples. ASR is short for attack success rate. Number of Images represents the number of experimental images to evaluate the attack performance. Stealthiness Attack means that the attack has a more naturalness score. Physical Attack denotes that the attack can be applied in the physical world. Transferable Attack represents that the attack has transferable attack ability. A lightweight model means that the model, that generates the adversarial patches, has low computational cost.
Evaluation and Analysis for Experimental Consequence: As Table V illustrates, both AdvPatch(Fig.5) and AdvTetxure(Fig.5) do not have much more attack performance on the YOLOv2 model in the physical world and the ASRs are 13.57% and 12.72%, respectively. The Natural Patch(Fig.5) has higher ASR and its value of ASR is 19.43%, but the other Natural Patch(Fig.5) does not perform as well as Fig.5. The ASR of the first Natural Patch(Fig.5) is almost close to the Benign Image(Fig.5). The MVPatches perform better than other adversarial patches in the physical world. The ASR of MVPatches are respectively 22.60% (Fig.5) and 26.33% (Fig.5). Consequently, we ensure the MVPatch has more attack performance than others in the physical world.
V-C2 Evaluation of Attack Performance from Diverse Angles in the Physical Domain
We collected 6846 images from various angles ($0^{\circ}$, $30^{\circ}$, $60^{\circ}$, $90^{\circ}$) to assess the impact of different angles on the attack performance of adversarial patches, including the Natural Patch, AdvPatch, AdvTexture, Benign Image, and our MVPatch. To keep the fairness of the experiment, we change angles from {$0^{\circ}$, $30^{\circ}$, $60^{\circ}$, $90^{\circ}$} with the same distance (2 meters). The volunteer rotated while the adversarial patch rotated in the opposite direction, ensuring that the patch always faced the camera. We refer to this setup as ”with Rotation”. The experimental results are presented in Fig.6(a).
Explanation for Experimental Consequence: The light blue dotted line with squares represents the MVPatch with Rotation, while the red dotted line with circles represents the Benign Image with Rotation. The dark blue dotted line with triangles represents the AdvPatch with Rotation, while the green dotted line with triangles represents the AdvTexture with Rotation. The black dotted line with squares represents the Natural Patch with Rotation. The X-axis represents the angles, while the Y-axis represents the attack success rate.
Evaluation and Analysis for Experimental Consequence: The effectiveness of adversarial patches with rotation diminishes as the angle of rotation increases. At an angle of $0^{\circ}$, the ASR of the MVPatch can reach 39.5%, and other patches perform well at the angle of $0^{\circ}$ except for the AdvTexture. At the angle of $90^{\circ}$, the ASRs of the Natural Patch and AdvTexture surpass the MVPatch at the same angle. Even a slight variation in the angle can significantly enhance the robustness and aggressiveness of the adversarial patch when it is rotated. Fig.6(a) also demonstrates that the MVPatch outperforms other attack methods in terms of angle variations (at the angles of $0^{\circ}$, $30^{\circ}$, and $60^{\circ}$).
V-C3 Evaluation of Attack Performance from Diverse Distance in the Physical Domain
A total of 3984 images are collected at various distances (1 meter, 2 meters, 3 meters, and 4 meters) to assess how different distances impact the attack performance of adversarial patches, specifically MVPatch, Natural Patch, AdvPatch, AdvTexture, and Benign Image. To keep the fairness of the experiment, we change distances from {1, 2, 3, 4} meters with the same angle ($0^{\circ}$). Fig.6(b) presents the results.
Explanation for Experimental Consequence: MVPatch is represented by the light blue dotted line with squares. Natural Patch is represented by the black dotted line with squares. AdvPatch is represented by the dark blue dotted line with triangles. AdvTexture is represented by the green dotted line with triangles. The benign Image is represented by the red dotted line with circles. The X-axis represents the attack success rate, and the Y-axis represents the distance from the volunteer with the adversarial patch to the camera.
Evaluation and Analysis for Experimental Consequence: As Fig.6(b) shows, the adversarial patches generated by MVPatch and AdvPatch exhibit high levels of aggression when the volunteer is at a distance of one meter from the camera. The ASR of MVPatch is 71.67% and the ASR of AdvPatch is 63.46% at a distance of 1 meter. As the volunteer moves farther away from the camera (beyond one meter), the effectiveness of the attack diminishes. However, the adversarial patches generated by AdvTexture and Natural Patch have higher aggression at a distance of 2 meters. When the volunteer moves farther away from the camera (beyond two meters), the effectiveness of the attack diminishes. Similarly, the attack weakens as the volunteer moves closer.
V-D Illustraton of Experimental Consequence
This section provides experimental results in both the digital and physical domains to demonstrate our proposed approach. The green rectangle represents a successful detection of a person by the object detector, while the red rectangle indicates a failure to detect the person class. In the digital domain, we choose various experimental images from the Inria Person dataset as evaluation objects. In the physical domain, the source images of this experiment are not covered with mosaics. For privacy protection, mosaics are later added near the face. Fig.7 displays the results of digital experiments, while Fig.8 depicts the results of physical experiments.
VI Conclusion
In this paper, we propose a dual-perception-based attack framework. This framework comprises a model-perception degradation method and a human-perception improvement method. The aim is to create vivid and aggressive adversarial patches, which we refer to as MVPatch, for person detection models suitable for real-world implementation. In terms of the model-perception degradation method, we exploit the ensemble attack on various object detectors to enhance the transferability of the adversarial patch. In terms of the human-perception improvement method, we introduce a lightweight approach for visual similarity measurement designed to make the adversarial patch less noticeable. Through extensive qualitative and quantitative experiments, we compare different approaches and find that minimizing the object loss yields the most effective patches while maintaining competitive attack performance compared to similar methods. We evaluate the transferability and naturalness of the crafted adversarial patches in both digital and physical domains. Extensive experiments demonstrate that the proposed MVPatch performs remarkable stealthiness and transferability with few computational costs.
However, there are still limitations in terms of factors such as angles and distance that can affect the performance of adversarial attacks. In future work, we aim to explore approaches to designing adversarial patches that can minimize the influence of angle and distance and maximize attack performance in both digital and physical domains.

### v3

@@
 Abstract
-Recent investigations demonstrate that adversarial patches can be utilized to manipulate the result of object detection models.
-However, the conspicuous patterns on these patches may draw more attention and raise suspicions among humans.
-Moreover, existing works have primarily focused on enhancing the efficacy of attacks in the physical domain, rather than seeking to optimize their stealth attributes and transferability potential.
-To address these issues, we introduce a dual-perception-based attack framework that generates an adversarial patch known as the More Vivid Patch (MVPatch).
-The framework consists of a model-perception degradation method and a human-perception improvement method.
-To derive the MVPatch, we formulate an iterative process that simultaneously constrains the efficacy of multiple object detectors and refines the visual correlation between the generated adversarial patch and a realistic image.
-Our method employs a model-perception-based approach that reduces the object confidence scores of several object detectors to boost the transferability of adversarial patches.
-Further, within the human-perception-based framework, we put forward a lightweight technique for visual similarity measurement that facilitates the development of inconspicuous and natural adversarial patches and eliminates the reliance on additional generative models.
-Additionally, we introduce the naturalness score and transferability score as metrics for an unbiased assessment of various adversarial patches’ natural appearance and transferability capacity.
-Extensive experiments demonstrate that the proposed MVPatch algorithm achieves superior attack transferability compared to similar algorithms in both digital and physical domains while also exhibiting a more natural appearance.
-These findings emphasize the remarkable stealthiness and transferability of the proposed MVPatch attack algorithm.
+Recent studies have shown that Adversarial Patches (APs) can effectively manipulate object detection models.
+However, the conspicuous patterns often associated with these patches tend to attract human attention, posing a significant challenge.
+Existing research has primarily focused on enhancing attack efficacy in the physical domain while often neglecting the optimization of stealthiness and transferability.
+Furthermore, applying APs in real-world scenarios faces major challenges related to transferability, stealthiness, and practicality.
+To address these challenges, we introduce generalization theory into the context of APs, enabling our iterative process to simultaneously enhance transferability and refine visual correlation with realistic images.
+We propose a Dual-Perception-Based Framework (DPBF) to generate the More Vivid Patch (MVPatch), which enhances transferability, stealthiness, and practicality.
+The DPBF integrates two key components: the Model-Perception-Based Module (MPBM) and the Human-Perception-Based Module (HPBM), along with regularization terms.
+The MPBM employs ensemble strategy to reduce object confidence scores across multiple detectors, thereby improving AP transferability with robust theoretical support.
+Concurrently, the HPBM introduces a lightweight method for achieving visual similarity, creating natural and inconspicuous adversarial patches without relying on additional generative models.
+The regularization terms further enhance the practicality of the generated APs in the physical domain.
+Additionally, we introduce naturalness and transferability scores to provide an unbiased assessment of APs.
+Extensive experimental validation demonstrates that MVPatch achieves superior transferability and a natural appearance in both digital and physical domains, underscoring its effectiveness and stealthiness.
 Index Terms:
@@
 Deep Neural Networks (DNNs) have achieved remarkable performance in manifold fields, such as computer vision [4, 5, 6, 7], natural language processing[8, 9], and automatic speech recognition [10, 11].
-However, due to their lack of interpretability, DNNs are vulnerable to Adversarial Examples (AEs)[12, 13, 14], which raises concerns about their reliability in security-critical applications such as face recognition[15, 16] and autonomous driving[17, 18].
-In general, AE attacks can be categorized into two types: Digital Attack (DA)[12, 13, 19, 20, 21, 22, 23, 24, 25, 26], which involves introducing digital perturbations into the input image to carry out the attack, and Physical Attack (PA)[27, 28, 29, 30, 31, 2, 32, 1, 33, 34, 3, 35, 16, 36, 37, 38], which directly targets real-world objects using adversarial perturbations, as illustrated in Fig.1(b) and Fig.1(a), respectively.
-Object Detection Models (ODMs)[39, 40, 41], comprised of deep neural layers, are extensively utilized in real-world applications such as people tracking[42], pedestrian and vehicle re-identification (Re-ID)[43, 44], and remote sensing[45].
+However, due to their lack of interpretability, DNNs are vulnerable to Adversarial Examples (AEs) [12, 13, 14, 15], which raises concerns about their reliability in security-critical applications such as face recognition [16, 17] and autonomous driving [18, 19].
+In general, AE attacks can be categorized into two types: Digital Attack (DA) [12, 13, 20, 21, 22, 23, 24, 25, 26, 27], which involves introducing digital perturbations into the input image to carry out the attack, and Physical Attack (PA) [28, 29, 30, 31, 32, 2, 33, 1, 34, 35, 3, 36, 17, 37, 38, 39], which directly targets real-world objects using adversarial perturbations, as illustrated in Fig.
+1(b) and Fig.
+1(a), respectively.
+Object Detection Models (ODMs) [40, 41, 42], comprised of deep neural layers, are extensively utilized in real-world applications such as people tracking [43], pedestrian re-identification (Re-ID) [44], and remote sensing [45].
 Although ODMs enhance human convenience and improve the quality of life, they simultaneously pose a considerable risk to individual privacy and sensitive personal information [1, 3, 2, 33].
-Furthermore, the capability of physical attacks to be transferred amplifies the vulnerability of ODMs[46, 47, 48].
+Furthermore, the capability of physical attacks to be transferred amplifies the vulnerability of ODMs [46, 47, 48, 49, 50, 51].
 In real-world scenarios, AEs could be often observed which are quasi-imperceptible to humans.
-For example, as depicted in Fig.1, an intruder appears under a surveillance camera holding an adversarial patch, which is camouflaged as Van Gogh’s painting, and the DNN-based object detection system will fail.
+For example, as depicted in Fig.
+1, an intruder appears under a surveillance camera holding an adversarial patch, which is camouflaged as Van Gogh’s painting, and the DNN-based object detection system will fail.
 It can even be said that the intruder is invisible under the camera, which will seriously threaten the public safety system Hence, it is imperative to conduct a thorough investigation into PA on ODMs.
 Adversarial Patches (APs) [31] have emerged as a potent means to manipulate ODMs within the physical realm, offering several advantages such as input-independence and scene-independence, with substantial real-world impacts [23, 16, 52, 53, 3, 34].
-However, much of the extant literature emphasizes enhancing the efficacy of the attack within this physical sphere, rather than seeking to optimize its stealth attributes and transferability potential.
-Hu et al.[1] considered how to improve the stealthiness of the adversarial patch by Generative Adversarial Network (GAN)[51].
-In contrast, Guesmi et al.[38] proposed a lightweight approach to generate natural patches without GAN.
-Moreover, many scholars have endeavored to augment APs’ transferability.
-A case in point is the research by Wang et al.[46], who honed AP transferability within automated checkout systems through a bias-guided arrangement.
-In a similar endeavor, Wei et al.[47] devised a Pay No Attention (PNA) attack, wherein attention gradients are disregarded during backpropagation, thereby effectively improving backpropagation’s linearity.
-Furthermore, Huang et al.[48] modified a single surrogate model into a diverse ensemble structure, yielding more transferable adversarial examples.
-Nevertheless, numerous challenges persist, which include the substantial computational resource requirements associated with the use of GAN during the training process, and an unrelenting research focus on enhancing the stealthiness[2, 31, 52, 1, 32, 3, 35, 29, 53, 28, 38] and transferability[46, 47, 48] of APs.
-Significantly, there remains limited research simultaneously addressing both stealthiness and transferability issues related to APs, with an additional focus on computational resource efficiency.
-To bridge this gap, we propose a dual-perception-based framework to generate a potent adversarial patch, referred to as the MVPatch (More Vivid Patch).
-This attack framework offers remarkable stealthiness and transferability capacities for attacking against ODMs.
-To improve the transferability potential, we present a model-perception degradation method for ensemble multiple object detectors achieved through the ensemble attack loss function.
-Additionally, to enhance the stealthiness attributes, we present a human-perception improvement method realized by the Compared Specified Image Similarity (CSS) loss function, which allows for the generation of natural and stealthy adversarial patches without the reliance on additional generative models.
-By integrating both the model-perception-based method and the human-perception-based method, we can generate a more powerful adversarial patch with significant transferable attack ability and invisibility.
-Moreover, we introduce the naturalness score and the transferability score as experimental evaluation metrics to assess the attack performance regarding stealthiness and transferability.
-We exploit the YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny as threat models and Faster RCNN, SSD, and YOLOv5 as transfer attack models.
-Based on extensive experiments conducted in both digital and physical scenarios, along with several independent subjective surveys, our proposed MVPatch consistently generates more natural and transferable results compared to several state-of-the-art baselines.
-We present the primary specialization of MVPatch, along with the distinguishing features of adversarial patches that possess meaning and those that are devoid of meaning.
-The main contributions of this work can be summarized as follows:
-  • To the best of our knowledge, we are the first to propose a dual-perception-based attack framework for generating an adversarial patch (MVPatch).
-This patch offers both stealthiness and transferability by exploiting both model-perception-based and human-perception-based methods.
-  • We present the model-perception-based method, an attack method that reduces the object confidence scores of an ensemble of multiple object detectors using an ensemble attack loss function.
-This enhances the transferability of the adversarial patches, significantly improving attack performance in both digital and physical domains compared to state-of-the-art baselines.
-  • We propose a human-perception-based method, a lightweight visual similarity measurement method that generates natural and stealthy adversarial patches without the need for additional generative models.
-This allows us to generate patches based on specific images.
-  • We employ naturalness and transferability scores as evaluation metrics to assess the naturalness and attack transferability of the diverse adversarial patches in our experiment.
+However, despite their potential, adversarial patches face three major issues that hinder their broader applicability and effectiveness: Transferability, Stealthiness, and Practicality.
+Transferability can be enhanced through methods like gradient-based optimization (e.g., MI [54], NI [55]), input transformation (e.g., DI [56], SI [55]), and model ensemble methods (e.g., MI [54], SVRE [57]), with the latter being the most effective yet not fully understood.
+Stealthiness involves generating realistic APs using GAN-based methods (e.g., Natural Patch [1]), Diffusion-based methods (e.g., Diff-PGD [38]), and Regularization-based methods (e.g., AdvART [39], DAP [58]), each with trade-offs in computational cost and efficiency.
+Practicality concerns the transition from digital to physical domains [30, 29, 28, 35], where APs face challenges like performance degradation due to varying distances and angles.
+Moreover, much of the existing literature either focuses on enhancing the transferability of adversarial patches [46, 47, 48, 59] while overlooking their stealthiness, or prioritizes stealthiness [2, 32, 60, 1, 33, 3, 36, 30, 61, 29, 39] at the expense of transferability.
+Notably, there is limited research that simultaneously addresses both stealthiness and transferability, while also considering computational resource efficiency.
+To address these challenges, we propose a Dual-Perception-Based Framework (DPBF) for generating a powerful adversarial patch, termed the More Vivid Patch (MVPatch).
+This framework enhances the transferability, stealthiness, and practicality of adversarial patches for attacking object detection models (ODMs).
+To improve transferability, we introduce the Model-Perception-Based Module (MPBM), which employs an ensemble strategy with multiple object detectors.
+The MPBM leverages the ensemble strategy to boost adversarial attack transferability, supported by theoretical insights from generalization theory.
+For enhanced stealthiness, we present the Human-Perception-Based Module (HPBM).
+This module generates adversarial patches that closely mimic a specified image in Hilbert space, using diverse transformations and the Frobenius norm to minimize generalization error.
+This approach ensures the patches are natural and inconspicuous without relying on additional generative models.
+To improve real-world practicality, we incorporate regularization terms such as Total Variation (TV) and Non-Printable Score (NPS) into the DPBF.
+By integrating the MPBM, HPBM, and regularization terms, we create a more effective adversarial patch with significant transferability and stealthiness while maintaining practicality.
+Additionally, we introduce naturalness and transferability scores as experimental metrics to assess the performance of the adversarial patches.
+We evaluate these patches using threat models including YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny, as well as transfer attack models such as Faster R-CNN, SSD, and YOLOv5.
+Extensive experiments in both digital and physical scenarios, along with independent subjective surveys, demonstrate that our MVPatch consistently outperforms several state-of-the-art baselines in terms of naturalness and transferability.
+The main contributions can be summarized as follows:
+  • To the best of our knowledge, we are the first to introduce generalization theory to the field of adversarial patches.
+This approach provides robust theoretical support for generating patches that maintain transferability, stealthiness, and practicality through our Dual-Perception-Based Framework (DPBF).
+  • We present the Model-Perception-Based Module (MPBM), which reduces object confidence scores through an ensemble strategy.
+Utilizing generalization theory, the MPBM demonstrates that this ensemble approach enhances both generalization and stability, thereby improving the transferability of adversarial patches (APs) compared to single-model strategies.
+  • We propose a Human-Perception-Based Module (HPBM) that generates natural and stealthy adversarial patches without the need for additional generative models, while ensuring minimal generalization error.
+  • We employ naturalness and transferability scores as evaluation metrics to assess the naturalness and attack transferability of diverse adversarial patches in our experiments.
   • We perform a comprehensive analysis of the transfer attack performance and naturalness of the proposed method, using both meaningful and meaningless patch approaches in a variety of digital and physical scenarios.
-The remainder of this paper is organized as follows.
-Section II provides a concise literature review on adversarial examples and adversarial patches in various scenarios, encompassing both digital and physical contexts.
-In Section III, we introduce the definition and characteristics of adversarial patch attacks, along with detailed information regarding the input and output of the threat models.
-The generation process of our proposed MVPatch is then presented in this section.
-Section IV introduces the environment and dataset in the experiment, followed by the provision of relevant experimental details.
-Additionally, we propose evaluation metrics for the experiment.
-Section V presents empirical evidence that showcases the effectiveness of the proposed MVPatch through comprehensive evaluations conducted in both digital and physical environments.
-Ultimately, we summarize the paper and present the further research directions in Section VI.
+The remainder of this paper is organized as follows: Section II reviews the literature on adversarial examples, covering both digital and physical domains, and explores generalization aspects for adversarial attacks.
+Section III defines adversarial patch attacks, discusses their characteristics, and details the input and output of the threat models, as well as the generation process of the proposed MVPatch.
+Section IV describes the experimental environment and dataset, including the experimental setup and evaluation metrics.
+Section V demonstrates the empirical effectiveness of MVPatch through extensive evaluations in both digital and physical settings.
+Finally, Section VI summarizes the findings, discusses limitations, and outlines future research directions.
 II Related Work
-In this section, we briefly review the literature related to adversarial examples and adversarial patches in diverse scenarios, such as in the digital domains and physical domains.
-Furthermore, we discuss the recent issues to be solved and present the necessity of our work.
 II-A Digital Attack
-Adversarial Examples (AEs) are elaborately designed perturbations that are imperceptible to humans but could mislead Deep Neural Networks (DNNs).
-In 2014, Szegedy et al.[12] discovered a special property of DNNs known as Adversarial Examples (AEs).
-They successfully attacked a DNN using the L-BFGS algorithm, causing it to produce incorrect predictions.
-In 2015, Goodfellow et al.[13] introduced the Fast Gradient Sign Method (FGSM), which improved the effectiveness and success rates of attacks on neural network classification models.
-Kurakin et al.[22] developed the Basic Iterative Method (BIM) and demonstrated the application of AEs in the physical world.
-Papernot et al.[23] focused on partial attack and presented the Jocobian-Based Saliency Map Attack (JSMA) algorithm.
-Madry et al.[19] proposed Project Gradient Descent (PGD), an innovative framework that combines attack and defense algorithms to enhance the robustness of DNNs.
-There are numerous varieties of adversarial attacks in the digital domain, e.g., Deepfool[21], C&W[20], ZOO[24], Universal Perturbation[25], One-Pixel attack[26] and some attacks on large language vision models[54, 55, 56].
+Adversarial Examples (AEs) are perturbations that are imperceptible to humans but can mislead Deep Neural Networks (DNNs).
+In 2014, Szegedy et al.
+[12] first discovered AEs, successfully attacking a DNN using the L-BFGS algorithm.
+In 2015, Goodfellow et al.
+[13] introduced the Fast Gradient Sign Method (FGSM), enhancing attack effectiveness on neural network classifiers.
+Kurakin et al.
+[23] developed the Basic Iterative Method (BIM) and demonstrated AEs’ applicability in the physical world.
+Papernot et al.
+[24] presented the Jocobian-Based Saliency Map Attack (JSMA) for partial attacks.
+Madry et al.
+[20] proposed Project Gradient Descent (PGD) to enhance DNN robustness through a combined attack and defense framework.
+Other notable digital adversarial attacks include Deepfool [22], C&W [21], ZOO [25], Universal Perturbation [26], One-Pixel attack [27], and attacks on hasing retrival [62] and large language vision models [63, 64, 65].
 II-B Physical Attack
-Recently, it has been observed that printed AEs can effectively deceive neural network models in the physical domain.
-Athalya et al.[27] developed robust three-dimensional AEs that can mitigate the impact of diverse angle transformations.
-Sharif et al.[28] successfully attacked facial recognition systems by constructing adversarial eyeglasses, while Komkov et al.[29] proposed the construction of an adversarial hat for the same purpose.
-In 2018, Brown et al.[30] introduced the concept of Adversarial Patch (AP).
-Liu et al.[31] proposed DPatch to deceive object detection models, while at around the same time, Thy et al.[2] introduced AdvPatch to mislead automated surveillance systems.
+Recently, it has been observed that printed Adversarial Examples (AEs) can effectively deceive neural network models in the physical domain.
+Athalya et al.
+[28] developed robust 3D AEs mitigating diverse angle transformations.
+Sharif et al.
+[29] and Komkov et al.
+[30] attacked facial recognition systems with adversarial eyeglasses and hats, respectively.
+Brown et al.
+[31] introduced the concept of Adversarial Patches (APs).
+Liu et al.
+[32] proposed DPatch for object detection models, and Thy et al.
+[2] introduced AdvPatch for misleading surveillance systems.
 Wu et al.
-[32] presented the AdvCloack to make humans invisible in the object detector.
-Hu et al.[1] designed a more natural patch based on the AdvPatch[2].
-Xu et al.[33] proposed to construct an adversarial T-shirt to mislead the object detection systems.
-Huang et al.[34] introduced the Transfer-based Self-ensemble Attack (T-SEA) for attacking object detectors using ensemble attacks, effectively improving the transferable attack performance.
-Hu et al.[3] proposed the Toroidal-Cropping-based Expandable Generative Attack (TC-EGA) algorithm, which applies adversarial textures to clothing to improve the stealthiness of the adversarial patch.
-Wang et al.[35] constructed the adversarial patch based on the attention mechanism for the car camouflage.
-Wei et al.[16] investigated methods to enhance the attack performance of adversarial patches by focusing on their spatial positioning.
-Liu et al.[36] proposed a Perceptual-Sensitive Generative Adversarial Network (PS-GAN) to enhance the visual fidelity and attacking performance for the adversarial patches.
-Xue et al.[37] presented the Diffusion-Based Projected Gradient Desent (DIff-PGD) to generate realistic adversarial examples by using diffusion models.
-Guesmi et al.[38] proposed AdvART to attack object detectors by a simple method without generative models.
-Despite the significant results achieved by prior research on adversarial patches in the physical world, there are still some issues to be addressed, such as the trade-off between transferability[34, 1, 57] and stealthiniess[27, 29, 32, 33, 3, 35, 58, 36, 37, 38] of the adversarial patch.
-Furthermore, most recent research focuses on generating meaningless adversarial patches.
-The meaningless adversarial patch means that has no sense for human eye perception as Fig.1(d) illustrates.
-To the best of our knowledge, only one study has systematically examined the investigation of meaningful and transferable adversarial patches through the utilization of a generative adversarial network (GAN) by Hu et al.[1].
-However, training a generative adversarial network is challenging and computationally expensive.
-To address these issues, we propose the MVPatch algorithm, which outperforms other state-of-the-art baseline methods in terms of transferable and stealthy attacks both in the digital and physical scenarios while maintaining a lightweight approach.
+[33] presented AdvCloak to make humans invisible to object detectors.
+Hu et al.
+[1] designed a more natural patch based on AdvPatch, and Xu et al.
+[34] created an adversarial T-shirt for object detection.
+Huang et al.
+[35] introduced T-SEA for ensemble attacks on object detectors.
+Hu et al.
+[3] proposed TC-EGA, applying adversarial textures to clothing for stealthiness.
+Wang et al.
+[36] developed an adversarial patch for car camouflage using attention mechanisms.
+Wei et al.
+[17] focused on enhancing attack performance via spatial positioning of patches.
+Liu et al.
+[37] proposed PS-GAN for improved visual fidelity and attack performance.
+Xue et al.
+[38] introduced DIff-PGD to generate realistic AEs using diffusion models.
+Guesmi et al.
+[39] proposed AdvART for attacking object detectors without generative models.
+II-C Generalization Analysis for Adversarial Patch Attack
+Generalization evaluates how well a model trained on a dataset performs on unseen data, typically using Empirical Risk Minimization (ERM) [66].
+Given data $\{(x_{i},y_{i}),i=1,\dots,n\}$ from a probability distribution $\mathcal{D}$ on $\Omega\times\{-1,1\}$, a class of functions $\mathcal{F}:\Omega\rightarrow\mathbb{R}$ and and a loss function $l$, ERM finds a minimizer of the empirical loss:
+$f^{*}=\operatorname*{arg\,min}_{f\in\mathcal{F}}\mathcal{L}_{\text{emp}}(f):=\operatorname*{arg\,min}_{f\in\mathcal{F}}\sum_{i}l(f(x_{i}),y_{i}).$ (1)
+The complexity of the model $\mathcal{F}$ can be assessed using approaches such as VC-dimension [67], Rademacher complexity [68], and covering numbers [69].
+These analyses typically provide bounds on the generalization gap, which is the difference between the empirical loss and the expected loss:
+$|\left[l(f^{*}(x),y)\right]-\mathcal{L}_{\text{emp}}(f^{*}(x))|<O^{*}(\sqrt{\frac{c}{n}}).$ (2)
+From Eq.
+2, we observe that the generalization error depends on both model complexity $c$ and the number of samples $n$.
+Recent studies have investigated the relationship between adversarial robustness and generalization [70, 71].
+Huang et al.
+[35] introduced T-SEA, which utilizes the insights from Eq.
+2 to improve the transferability of generated adversarial patches.
+Despite significant advancements in adversarial patches for the physical world, challenges persist, particularly in balancing transferability [35, 1, 72] and stealthiness [28, 30, 33, 34, 3, 36, 73, 37, 38, 39].
+Recent research often focuses on generating meaningless adversarial patches that lack visual coherence for human perception, as shown in Fig.
+1(d).
+Hu et al.
+[1] systematically explored meaningful and transferable adversarial patches using GANs, although training GANs is computationally expensive.
+Moreover, there is a lack of robust theoretical support for addressing the issues of transferability, stealthiness, and practicality.
+To address these challenges, we propose a Dual-Perception-Based Framework (DPBF) to generate the More Vivid Patch (MVPatch) with enhanced transferability, stealthiness, and practicality.
+Inspired by T-SEA [35], we incorporate generalization theory to comprehensively analyze and improve the generalization and stability of adversarial patches, enhancing both their transferability and stealthiness.
+Our MVPatch outperforms state-of-the-art methods in both digital and physical scenarios, achieving superior transferability and stealthiness with a lightweight approach.
 III Methodology
@@
 Algorithm 1 illustrates the complete procedure for the MVPatch algorithm.
-In this section, we first introduce the definition and characteristics of adversarial patch attacks and then detail the input and output information of threat models.
-Finally, we summarize the entire generation process of MVPatch in this section.
-III-A Problem Formulation
-The objective of this study is to generate a malicious image $\mathcal{I^{*}}$ with an adversarial patch from a benign image $\mathcal{I}$, to deceive object detection models.
-The complete process of generating the adversarial patch is illustrated in Fig.2.
-Typically, the malicious images with adversarial patches can be formulated as:
-*$\mathcal{I^{*}}=(1-\mathcal{M})\odot\mathcal{I}+\mathcal{M}\odot\mathcal{P},$ (1)
-where $\mathcal{M}$ represents the patch mask, which determines the size, shape, and location of the adversarial patch $\mathcal{P}$.
-The $\odot$ symbol denotes the Hadamard product, which multiplies the elements of the corresponding positions in the matrices $\mathcal{M}$, $\mathcal{I}$, and $\mathcal{P}$.
-We denote the object detection model as $f(x):x\in\mathbb{R}^{\mathbf{H\times W\times C}}\mapsto\mathcal{Y}\in\mathbb{R}^{6}$.
-$x$ is input image and $\mathcal{Y}$ is label of object.
-The generation of adversarial patches can be viewed as a constrained optimization problem:
-$f(\mathcal{I^{*}})\neq f(\mathcal{I})\quad\mathrm{s.t.}\,\lVert\mathcal{I^{*}}-\mathcal{I}\rVert\longrightarrow\mathcal{P},$ (2)
-where $\mathcal{P}$ represents the adversarial patch generated under the constraints of *$\lVert\mathcal{I^{*}}-\mathcal{I}\rVert\longrightarrow\mathcal{P}$, while satisfying $f(\mathcal{I^{*}})\neq f(\mathcal{I})$.
-Finding a closed-form solution for this optimization problem is not feasible due to the non-convex nature of the deep neural network model $f(\cdot)$.
-Therefore, to approximately solve this optimization problem, we utilize the following equation:
-$\underset{\mathcal{P}}{\mathrm{argmax}}\,\mathcal{L}(f(\mathcal{I^{*}}),f(\mathcal{I})),$ (3)
-where $\mathcal{L}$ represents the cross-entropy loss function.
-To solve the optimization problem, we can utilize the Adam optimizer, which is one of the most common and widely used optimization techniques.
-III-B Threat Models
-We apply the YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny models to generate adversarial patches.
-These patches are then evaluated on the YOLOv5, Faster-RCNN, and SSD models to assess their transferability in attacks.
-In the object context of the YOLO series, given the benign image $\mathcal{I}$, the objective of the adversarial attack is to jeopardize the object detector $f_{\theta}(x,\mathcal{Y})$, where $x$ belongs to $\mathcal{I}$, $\mathcal{Y}=\{\mathrm{a},\mathrm{b},\mathrm{w},\mathrm{h},\mathrm{obj},\mathrm{cls}\}$ and $\theta$ denotes the parameter of the object detector.
-We suppress the object confidence score by utilizing the adversarial patches.
-$\mathrm{a}$ and $\mathrm{b}$ are represented by the value of the horizontal and vertical coordinates of the detection frame.
-$\mathrm{w}$ and $\mathrm{h}$ are the width and height of the detection frame.
-$\mathrm{obj}$ and $\mathrm{cls}$ are object confidence scores and classification scores.
-Our focus is to suppress the person’s object confidence score by optimizing the use of adversarial patches.
-III-C Robustness of Adversarial Patch in the Physical Domain
-Previous studies have primarily focused on applying adversarial patches in the digital domain.
-However, it is important to recognize that adversarial patches can also appear in the physical domain.
-To enhance the effectiveness of adversarial patches in physical attacks, we propose the utilization of the Non-Printable Score (NPS)[28] and Total-Variation (TV)[3] as loss functions.
-These will aid in the creation of more aggressive and robust adversarial patches specifically designed for physical scenarios.
-We present the NPS for improving the robustness when we generate adversarial patches in the physical domain using digital patterns and to ensure compatibility with the iPad color gamut and printer gamut.
-The formulation of the NPS is as follows:
-$\mathcal{L_{\mathrm{NPS}}}=\sum_{c\in\mathcal{C},p\in\mathcal{P}}|p-c|,$ (4)
-where $\mathcal{P}$ and $\mathcal{C}$ represent the pixel values of the adversarial patch and the colors that the iPad can display, respectively.
-Both $p$ and $c$ are elements of $\mathcal{P}$ and $\mathcal{C}$.
-To improve the physical robustness of adversarial patches, we utilize the Euclidean distance ||$\cdot$|| ⋅ | to constrain the generation of adversarial patches.
-We also introduce the TV term, which enhances the smoothness of neighboring pixels. The TV formulation is defined as:
+In this section, we provide a comprehensive overview of our Dual-Perception-Based Framework (DPBF) for generating More Vivid Patch (MVPatch) applicable to both digital and physical domains.
+We start by outlining the motivation behind the development of MVPatch.
+Following this, we introduce the foundational concepts, including notations, definitions, and problem formulation.
+We then detail the specifics of our proposed adversarial patch attack framework.
+Finally, we summarize the entire MVPatch generation process.
+III-A Motivation
+Adversarial Patches (APs) have gained significant attention in adversarial machine learning for their ability to manipulate object detection models.
+These patches can be strategically placed to cause misclassifications, posing challenges for robust detection systems.
+However, APs face three major challenges: Transferability, Stealthiness, and Practicality.
+  • Transferability: There is insufficient theoretical support explaining why ensemble methods improve transferability, indicating a need for further research.
+  • Stealthiness: GAN-based and diffusion-based methods generate realistic APs but at high computational costs and longer processing times, presenting a trade-off between realism and efficiency.
+  • Practicality: In the physical domain, AP performance degrades with varying distances and angles, necessitating solutions for effective real-world deployment.
+To address these challenges, we propose a Dual-Perception-Based Framework (DPBF) designed to generate effective and stealthy adversarial patches, termed MVPatch.
+This framework improves the transferability of adversarial patches across various models while also enhancing their visual inconspicuousness and robustness in physical environments.
+We frame the adversarial transferability problem as an optimal generalization issue and introduce two modules to minimize generalization error: the Model-Perception-Based Module (MPBM), which enhances transferability through ensemble strategy, and the Human-Perception-Based Module (HPBM), which improves patch stealthiness using the Frobenius norm in Hilbert space transformations.
+Additionally, we incorporate regularization terms such as Total Variation (TV) and Non-Printable Score (NPS) into the DPBF to improve practical deployment in real-world scenarios.
+The overall framework is illustrated in Figure 2.
+III-B Preliminary
+Notations.
+We consider neural networks $\mathcal{F}(\mathcal{I},\mathcal{Y})$ for object detection tasks and let $\mathcal{I}$ be the input space of the model, and let $\mathcal{Y}=\{\mathrm{a},\mathrm{b},\mathrm{w},\mathrm{h},\mathrm{obj},\mathrm{cls}\}$.
+Here, $\mathrm{a}$ and $\mathrm{b}$ represent the horizontal and vertical coordinates of the detection frame, respectively.
+$\mathrm{w}$ and $\mathrm{h}$ denote the width and height of the detection frame.
+$\mathrm{obj}$ and $\mathrm{cls}$ are the object confidence scores and classification scores.
+We model the neural network as a mapping function $\mathcal{F}:\mathcal{I}\subseteq\mathbb{R}^{\mathbf{H\times W\times C}}\rightarrow\mathcal{Y}\subseteq\mathbb{R}^{N}$.
+The notations $f\in\mathcal{F}$, $x\in\mathcal{I}$, and $y\in\mathcal{Y}$ are the realizations of random variables.
+For a given input $x\in\mathcal{I}$, the object detection model $f\in\mathcal{F}$ first predicts the confidence score for each label $y_{\mathrm{obj}}\in\mathcal{Y}$, denoted as $f_{\mathrm{obj}}(x)$.
+These confidence scores sum up to 1, i.e., $\sum_{y_{\mathrm{obj}}\in\mathcal{Y}}f_{\mathrm{obj}}(x)=1,\forall x\in\mathcal{I}$.
+For a given model $f\in\mathcal{F}$, there is usually a model-dependent loss function $\mathcal{L}_{f\in\mathcal{F}}:x\in\mathcal{I}\times y\in\mathcal{Y}\rightarrow\mathbb{R}^{+}$, which is the composition of a differentiable training loss, (e.g., cross-entropy loss) $\mathcal{L}$ and model’s confidence score $f_{\mathrm{obj}}(\cdot)$.
+Thus, $\mathcal{L}(x,y):=\mathcal{L}(f(x),y)$, $(x,y)\in(\mathcal{I},\mathcal{Y})$.
+Definition 1 (Adversarial Patch)
+An adversarial patch is a perturbation $\mathcal{P}$ applied to a benign image $\mathcal{I}$ to produce a malicious image $\mathcal{I}^{*}$, with the intent of deceiving object detection models $f\in\mathcal{F}$.
+Formally, it can be expressed as:
+$f(\mathcal{I}^{)}\neq f(\mathcal{I})=\mathcal{Y},\quad\text{where}\quad\mathcal{I}^{*}=(1-\mathcal{M})\odot\mathcal{I}+\mathcal{M}\odot\mathcal{P},$ (3)
+where $\mathcal{M}$ denotes the patch mask that determines the size, shape, and placement of the adversarial patch $\mathcal{P}$.
+The symbol $\odot$ represents the Hadamard product, which involves element-wise multiplication of the corresponding entries in the matrices $\mathcal{M}$, $\mathcal{I}$, and $\mathcal{P}$.
+Definition 2 (Generalization Error)
+Consider a training dataset $\mathcal{T}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{N},y_{N})\}$ drawn from a joint probability distribution $P(\mathcal{I},\mathcal{Y})$.
+Let the hypothesis space consist of a set of models =$\mathcal{F}=\{f_{1},f_{2},\ldots,f_{M}\}$, where $M$ denotes the number of models.
+The generalization error $\mathcal{G}(f)$ for a model $f\in\mathcal{F}$ is defined as:
+$\mathcal{G}(f)\triangleq\underbrace{\mathbb{E}\left[\mathbbm{1}\{f(x)\neq y\}\right]}_{\text{Expected Error}}-\underbrace{\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}\{f(x_{i})\neq y_{i}\}}_{\text{Empirical Error}},$ (4)
+where $\mathbbm{1}\{\cdot\}$ is the indicator function that evaluates to 1 if the condition is true and 0 otherwise.
+Problem Formulation.
+To generate adversarial patches that are both transferable and stealthy, we introduce the generalization error $\mathcal{G}(\cdot)$ and formulate the problem as an optimization task.
+The objective is to find an image patch $p\in\mathcal{P}$ of dimensions $(H,W,C)$, along with transformations $t\in T$, that minimize the generalization error across multiple neural networks $f\in\mathcal{F}$:
+$\mathcal{P}=\operatorname*{arg\,min}_{\mathcal{P}\in\mathbb{R}^{H\times W\times C}}\mathcal{G}(f,x,p,t),\quad\text{s.t.}\quad d(t(p),t(s)),$ (5)
+where $d(\cdot)$ is the distance function that quantifies the similarity between the adversarial patch $p\in\mathcal{P}$ and a specified image $s\in\mathcal{S}$.
+Minimizing $\mathcal{G}(f,x,p,t)$ ensures that the generated adversarial patch maintains high transferability, while satisfying $d(t(p),t(s))$ guarantees its high stealthiness.
+III-C Model-Perception-Based Module
+To enhance the transferability of Adversarial Patches (APs), we propose the Model-Perception-Based Module (MPBM).
+The MPBM incorporates generalization error analysis to derive the generalization error bounds for our method.
+By comparing these bounds with those of single-model and ensemble-model strategies, we divide the ensemble approach into two distinct objective functions to validate the efficacy of our method.
+This enables us to substantiate our approach through both theoretical derivation and empirical verification.
+We begin with theoretical analyses focused on generalization and stability, identifying key factors that contribute to improved transferability across various models.
+Definition 3 (Ensemble Model)
+Given a set of models $\{f_{i}\}^{M}_{i=1}$, ensemble strategies can be divided into averaging outputs of neural networks and maximizing outputs of neural networks.
+By ensembling these models, we obtain a new predictive model:
+$f_{\text{ens}}(x)=\begin{cases}\frac{1}{M}\sum_{i=1}^{M}f_{i}(x)&f_{\text{ens}}^{\text{ave}}\,\,\,\text{for averaging outputs},\\ \max\{f_{i}(x)\}^{M}_{i=1}&f_{\text{ens}}^{\text{max}}\,\,\,\text{for maximizing outputs}.\end{cases}$ (6)
+Remark 1
+Here, $f_{\text{ens}}^{\text{max}}$ is a special case where the ensemble model strategy can also be considered a separate model.
+Therefore, we analyze it as an independent model in a broad sense.
+All subsequent references to $f_{\text{ens}}$ refer specifically to $f_{\text{ens}}^{\text{ave}}$.
+III-C1 Generalization Analysis
+Theorem 1 (Generalization Error of Ensemble Model)
+Given an ensemble model $f_{\text{ens}}$ and a single model $f$ from the model set $\mathcal{F}$, where $(x,y)$ are drawn from the distribution $\mathcal{D}$, with $x$ and $y$ representing the input and output spaces of the models, respectively.
+Then,
+$\mathbb{E}_{(x,y)\sim\mathcal{D};f\sim\mathcal{F}}\left[\mathcal{L}(f_{\text{ens}}(x),y)\right]\leq\mathbb{E}_{(x,y)\sim\mathcal{D};f\sim\mathcal{F}}\left[\mathcal{L}(f(x),y)\right].$ (7)
+Proof:
+$\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathcal{L}(f_{\text{ens}}(x),y)\right]=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathcal{L}\left(\frac{1}{M}\sum_{i=1}^{M}f_{i}(x),y\right)\right]$ (8)
+[…]
+$=\mathbb{E}_{(x,y)\sim\mathcal{D};f\sim\mathcal{F}}\left[\mathcal{L}(f(x),y)\right].$ (10)
+Remark 2
+Since the loss function $\mathcal{L}$ is typically convex, we can utilize Jensen’s inequality to establish the relationship between the ensemble model $f_{\text{ens}}$ and single model $f$ as shown in Theorem 1.
+Specifically, Jensen’s inequality ensures that the loss of the ensemble model, which is a convex combination of the individual models, will not exceed the average loss of the individual models.
+This highlights that ensemble model generally exhibit better generalization error compared to individual models.
+III-C2 Stability Analysis
+Theorem 2 (Stability of the Ensemble Model)
+Let Var denote variance and Cov denote covariance.
+Suppose the individual neural networks $f_{i},f_{j}\in\mathcal{F}$ each have a variance of $\sigma^{2}$.
+The variance of the ensemble model $f_{\text{ens}}$ can be expressed as:
+$\text{Var}(f_{\text{ens}})=\frac{\sigma^{2}}{M}\left(1+2\rho\right),$ (11)
+where $\rho=\frac{\text{Cov}(f_{i}(x),f_{j}(x))}{\sqrt{\text{Var}(f_{i}(x))\text{Var}(f_{j}(x))}}$ is the correlation coefficient between pairs of models, and $M$ is the number of models.
+Proof:
+$\text{Var}(f_{\text{ens}})$ $=\text{Var}\left[\frac{1}{M}\sum_{i=1}^{M}f_{i}(x)\right]$ (12)
+$=\frac{1}{M^{2}}\text{Var}\left[\sum_{i=1}^{M}f_{i}(x)\right]$ (13)
+$=\frac{1}{M^{2}}\underbrace{\left[\sum_{i=1}^{M}\text{Var}\left(f_{i}(x)\right)+2\sum_{i\neq j}\text{Cov}\left(f_{i}(x),f_{j}(x)\right)\right]}_{\text{Bienaymé's identity}}$ (14)
+$=\frac{1}{M^{2}}\left(M\sigma^{2}+2M\rho\sigma^{2}\right)$ (15)
+$=\frac{\sigma^{2}}{M}\left(1+2\rho\right).$ (16)
+Remark 3
+The variance of the ensemble model $f_{\text{ens}}$ is proportional to the correlation coefficient $rho$ and inversely proportional to the number of models $M$.
+Consequently, Eq.
+11 suggests that as the number of models $M$ increases, the variance will decrease, provided that the correlation between the models $\rho$ remains low.
+This reduction in variance implies an enhancement in both the stability and the generalization capability of the ensemble model.
+Thus, the adversarial patch generation process $\hat{x}\in\mathcal{I^{*}}$ using the ensemble model $f_{\text{ens}}$ can achieve greater stability compared to a single model approach, including strategy that aims to maximize outputs in an ensemble model.
+Therefore, we adopt the ensemble model strategy as outlined in Definition 3. The corresponding objective function is formulated as:
+$\mathcal{L}_{\text{OBJ}}=\mathbb{E}_{f\sim\mathcal{F}}\mathbb{E}_{\hat{x}\sim\mathcal{I^{*}}}f_{\text{ens}}^{\text{ave}}(\hat{x}).$ (17)
+The MPBM leverages the averaged outputs of the ensemble model $f_{\text{ens}}^{\text{ave}}$, ensuring enhanced stability and generalization during the adversarial patch generation process.
+By incorporating multiple models with low correlation, we achieve a more robust and reliable performance, leading to the conclusion that the ensemble strategy effectively mitigates variance and improves overall model efficacy.
+III-D Human-Perception-Based Module
+To enhance the stealthiness of Adversarial Patches (APs), we present the Human-Perception-Based Module (HPBM).
+This approach employs Kullback-Leibler (KL) divergence as a distance function and introduces a novel metric, the Compared Specified Image Similarity (CSS) measurement, to assess the similarity between the generated patch and a reference image.
+This methodology is analogous to the regularization techniques used in AdvART [39] and DAP [58].
+While we initially considered ensemble methods, their higher computational costs compared to single models led us to adopt a more streamlined approach.
+Consequently, we chose a lightweight method that preserves the realism and stealthiness of APs while avoiding the computational overhead associated with ensemble strategy.
+Definition 4 (Visual Reality of Adversarial Patches)
+Let $\mathcal{P}$ denote the set of image patches and $\mathcal{S}$ represent the set of specified reference images.
+To ensure visual realism, we aim to minimize the Kullback-Leibler (KL) divergence between $\mathcal{P}$ and $\mathcal{S}$.
+By employing Pinsker’s inequality, the lower bound of the KL divergence can be expressed as:
+$\mathcal{P}=\arg\min\mathbb{E}_{\hat{x}\sim\mathcal{P}}\left[\mathbb{KL}(\mathcal{P}||\mathcal{S})\right]\geq\frac{1}{2}\mathbb{E}_{\hat{x}\sim\mathcal{P}}\|\mathcal{P}-\mathcal{S}\|^{2}.$ (18)
+Theorem 3 (Upper Bound of Generalization Error)
+Consider a lower threshold of generalization error $t$ where $t\rightarrow 0$ and $t>0$.
+As defined in Definition 2, let $\mathcal{G}(f)$ denote the generalization error of an ensemble of $M$ models evaluated on $N$ data points.
+The generalization error $\mathcal{G}(f)$ can be bounded with a confidence level of $1-\gamma$, where $\gamma\in(0,1)$, as follows [35, 74]:
+$\mathcal{G}(f)\leq t(M,N,\gamma)=\sqrt{\frac{1}{2N}(\log M+\log\frac{1}{\gamma})}$ (19)
+Under Theorem 3, we augment the dataset with additional transformation operations to enhance data generalization.
+We employ the Frobenius norm in Hilbert space to design stealthy adversarial attacks, introducing the Compared Specified Image Similarity (CSS) measurement.
+The CSS metric is defined as follows:
+$\mathcal{L}_{\text{CSS}}=\mathbb{E}_{p\sim\mathcal{P}}\mathbb{E}_{t\sim\mathcal{T}}\|t(p)-t(s)\|_{\mathcal{F}}^{2},$ (20)
+where $t\in\mathcal{T}$ represents transformation operations such as rotation, flipping, and cropping.
+Intuition: There exists an inherent trade-off between the naturalness and the attack efficacy of adversarial patches.
+Higher attack performance generally results in lower naturalness, manifesting as a reduced resemblance to the source image.
+Conversely, increased naturalness often corresponds to diminished attack success rates.
+Further details are provided in Section V-B.
+III-E Practicality of Adversarial Patches in the Physical Domain
+Previous research has predominantly concentrated on the application of Adversarial Patches (APs) in the digital domain.
+However, it is essential to recognize that APs can also be effectively utilized in physical settings.
+To enhance the practicality of APs for physical environments, we propose incorporating Total Variation (TV) [3] and Non-Printable Score (NPS) [29] as regularization terms.
+These terms facilitate the creation of more aggressive and robust adversarial patches, specifically tailored to perform effectively in real-world physical environments.
+The Total Variation (TV) term is introduced to promote the smoothness of neighboring pixels. The TV regularization is defined as follows:
 $\mathcal{L_{\mathrm{TV}}}=\sum_{a,b}|p_{a,b}-p_{a,b+1}|+|p_{a,b}-p_{a+1,b}|,$ (21)
-where $p$ denotes the pixel values of the image, while $a$ and $b$ represent the abscissa and ordinate values of a single pixel, respectively.
-III-D Human-perception-based Method
-There is a trade-off between the naturalness and attack performance of adversarial patches.
-For instance, if an adversarial patch exhibits higher attack performance, it tends to have lower naturalness.
-In other words, as the similarity to the source image decreases, the success rate of the attack increases.
-To obtain a more natural adversarial patch, we propose the human-perception-based method by utilizing the Minkowski distance as a loss function to constrain the naturalness of the patch, which we refer to as the Compared Specified Image Similarity (CSS) measurement.
-The CSS is formulated as follows:
-$\mathcal{L_{\mathrm{CSS}}}=\left(\sum_{i=1}^{N}|\hat{x_{i}}-x_{i}|^{p}\right)^{\frac{1}{p}},$ (6)
-where $\hat{x}$ represents the adversarial patch generated using the specified image $x$. We choose $p=2$ in this work and $i$ represents the number of pixels of $\hat{x}_{i}$ and $x_{i}$.
-III-E Model-perception-based Method
-To improve the transferability of adversarial patches, we propose the model-perception-based method.
-First, we present the use of two distinct loss functions to attack object detection models, as shown below:
-$\mathcal{L}_{\mathrm{OBJ_{1}}}=\frac{1}{N}\sum_{i=1}^{N}S^{i}_{\mathrm{OBJ}}(\mathcal{I^{*}},\theta),$ (7)
-$\mathcal{L}_{\mathrm{OBJ_{2}}}=\mathop{\max}_{\mathcal{I^{*}}}(\{S^{i}_{\mathrm{OBJ}}(\mathcal{I^{*}},\theta)\}^{N}_{i=1}),$ (8)
-where $S^{i}_{\mathrm{OBJ}}$ denotes the object confidence score of the $i^{th}$ object detection model in the ensemble of object detection models (YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny).
-$\mathcal{I^{*}}$ is the image with adversarial patch and $\theta$ corresponds to the parameters of the object detection model.
-TABLE I: Attack Performance and Comprehensive Evaluation with Diverse Objective Loss Functions
-Faster-RCNN SSD YOLOv2 YOLOv3 YOLOv3-tiny YOLOv4 YOLOv4-tiny YOLOv5 Mean mAP CS ED
-$\mathcal{L}_{\text{OBJ}_{1}}$ 42.07% 54.57% 30.26% 31.39% 13.49% 27.31% 22.74% 35.72% 32.20% 97.91% 0.0428
-$\mathcal{L}_{\text{OBJ}_{2}}$ 49.40% 61.08% 50.34% 37.47% 41.44% 52.64% 56.46% 56.46% 51.80% 97.91% 0.0419
-We compare the attack performance of diverse object confidence score loss functions on various object detection models.
-It is observed that the adversarial patch generated by $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ outperforms the one generated by $\mathcal{L}_{\mathrm{{OBJ}_{2}}}$ in terms of attacking the object detection model.
-The adversarial patches generated through diverse loss functions are then applied to eight object detection models (Faster-RCNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5 models) to assess their attack performance and similarity to specified images.
-The adversarial patch generated by $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ not only exhibits excellent attack capability (the average mAP of $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ is 32.20% and the average mAP of $\mathcal{L}_{\mathrm{{OBJ}_{2}}}$ is 51.80%) but also maintains a high level of naturalness.
-Consequently, we select $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ as our objective loss function.
-The results of the experiment are illustrated in Table I, where the Faster-RCNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5 models are included as threat models.
-The mean mAP represents the average value of the mean Average Precision (mAP) obtained from diverse models.
-CS corresponds to the Cosine Similarity and ED denotes the Euclidean Distance.
-III-F Dual-perception-based Framework to Generate MVPatch
-To make the optimization process converge faster, we introduce additional optimization factors to the previously mentioned loss functions and formulate an ensemble loss function as Equation (9):
+where $p\in\mathcal{P}$ denotes the adversarial patch, and the indices $a$ and $b$ represent the horizontal and vertical coordinates, respectively.
+Additionally, we incorporate the Non-Printable Score (NPS) to enhance the robustness of adversarial patches when transitioning from digital to physical domains, ensuring compatibility with both the iPad color gamut and printer gamut.
+The NPS is formulated as follows:
+$\mathcal{L_{\mathrm{NPS}}}=\sum_{a,b}|p_{a,b}-c_{a,b}|,$ (22)
+where $p\in\mathcal{P}$ denotes the adversarial patch and $c\in\mathcal{C}$ represents the color matrix corresponding to the display capabilities of the iPad.
+The indices $a$ and $b$ refer to the horizontal and vertical coordinates, respectively.
+To further enhance the physical robustness of the adversarial patches, we employ the Euclidean distance ||$\cdot$|| ⋅ | as a constraint in the generation process.
+III-F Dual-Perception-Based Framework
+The Dual-Perception-Based Framework (DPBF) integrates the Model-Perception-Based Module (MPBM) and the Human-Perception-Based Module (HPBM) through the incorporation of Total Variation (TV), Non-Printability Score (NPS), and other hyperparameters, addressing the issues delineated in the Motivation section.
+The MPBM enhances the transferability of adversarial patches, while the HPBM ensures these patches are both realistic and stealthy.
+Furthermore, the inclusion of NPS and TV terms significantly augments the practicality of adversarial patches (APs), facilitating their seamless application from the digital domain to physical environments.
+To expedite the optimization process, additional optimization factors have been introduced into the existing loss functions, culminating in the formulation of an aggregate loss function as presented in Eq.
+(23):
 $\mathcal{L_{\mathrm{TOTAL}}}=\alpha\mathcal{L_{\mathrm{NPS}}}+\beta\mathcal{L_{\mathrm{TV}}}+\lambda\mathcal{L_{\mathrm{CSS}}}+\mathcal{L_{\mathrm{OBJ}}}.$ (23)
-Since optimizing the ensemble loss function with the Adam optimizer is challenging, we modified our optimization strategy.
-Instead of reducing the learning rate as training epochs increase, we set thresholds for the parameters and decrease the learning rate when a parameter no longer exhibits optimization during the training process.
-The updating strategy for the learning rate is governed by Equation (10):
-$\mathcal{LR_{\mathrm{N}}}=\gamma\mathcal{LR_{\mathrm{O}}},$ (10)
-where $\mathcal{LR}_{\mathrm{N}}$ is the new learning rate, while $\mathcal{LR}_{\mathrm{O}}$ corresponds to the old learning rate.
-In our experiments, we utilize $\gamma=0.01$ as the decay coefficient.
-The complete process is summarized in Algorithm 1.
-This algorithm aims to generate an adversarial patch, referred to as MVPatch, for a given original image $\mathcal{I}$ using a neural network $f(x)$ with the output being the adversarial patch $\mathcal{P}$.
 Algorithm 1 Generating Adversarial Patch (MVPatch)
@@
 4:  Initialize patch $\mathcal{P}$ using $\mathcal{M}$
-5:  *$\mathcal{I^{*}}=(1-\mathcal{M})\odot\mathcal{I}+\mathcal{M}\odot\mathcal{P}$
+5:  $\mathcal{I^{*}}=(1-\mathcal{M})\odot\mathcal{I}+\mathcal{M}\odot\mathcal{P}$
 6:  $\mathcal{L}_{\mathrm{\text{TOTAL}}}\leftarrow\alpha\mathcal{L}_{\mathrm{\text{NPS}}}+\beta\mathcal{L}_{\mathrm{\text{TV}}}+\lambda\mathcal{L}_{\mathrm{\text{CSS}}}+\mathcal{L}_{\mathrm{\text{OBJ}}}$
@@
 17:  Return: Adversarial Patch $P$
+Since optimizing the ensemble loss function with the Adam optimizer is challenging, we modified our optimization strategy.
+Instead of reducing the learning rate as training epochs increase, we set thresholds for the parameters and decrease the learning rate when a parameter no longer exhibits optimization during the training process.
+The updating strategy for the learning rate is governed by Eq.
+(24):
+$\mathcal{LR_{\mathrm{N}}}=\gamma\mathcal{LR_{\mathrm{O}}},$ (24)
+where $\mathcal{LR}_{\mathrm{N}}$ is the new learning rate, while $\mathcal{LR}_{\mathrm{O}}$ corresponds to the old learning rate.
+In our experiments, we utilize $\gamma=0.01$ as the decay coefficient.
+The complete process is summarized in Algorithm 1.
+This algorithm aims to generate an adversarial patch, referred to as MVPatch, for a given original image $\mathcal{I}$ using a neural network $f(x)$ with the output being the adversarial patch $\mathcal{P}$.
 IV Experiments
-In this section, we briefly introduce the environment and dataset in the experiment.
-Additionally, we propose evaluation metrics for the experiment.
-In the digital environment, we generate meaningful and meaningless adversarial patches and compare them with similar patch attack approaches.
-In the physical environment, we compare the Attack Success Rate (ASR) of MVPatch with similar patch attack methods and vary the angles and distances between the patch and camera to investigate their impacts on the ASR.
+In this section, we provide a brief introduction to the environment and dataset used in the experiment.
+Additionally, we delineate the threat models considered in the experiment and introduce the evaluation metrics employed.
+In the digital environment, we generate both meaningful and meaningless adversarial patches and compare them with similar patch attack approaches.
+In the physical environment, we compare the Attack Success Rate (ASR) of MVPatch with similar patch attack methods and vary the angles and distances between the patch and the camera to investigate their impacts on the ASR.
 IV-A Implementation Details
@@
 For collecting the source images, we utilize the iPhone 11 camera, which serves as our image input device.
-IV-B Experimental Environment and Dataset
+IV-B Threat Models
+We apply the YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny models to generate adversarial patches.
+These patches are then evaluated on the YOLOv5, Faster-RCNN, and SSD models to assess their transferability in attacks.
+In the object context of the YOLO series, given the benign image $\mathcal{I}$, the objective of the adversarial attack is to jeopardize the object detector $f_{\theta}(x,\mathcal{Y})$, where $x$ belongs to $\mathcal{I}$, $\mathcal{Y}=\{\mathrm{a},\mathrm{b},\mathrm{w},\mathrm{h},\mathrm{obj},\mathrm{cls}\}$ and $\theta$ denotes the parameter of the object detector.
+We suppress the object confidence score by utilizing the adversarial patches.
+$\mathrm{a}$ and $\mathrm{b}$ are represented by the value of the horizontal and vertical coordinates of the detection frame.
+$\mathrm{w}$ and $\mathrm{h}$ are the width and height of the detection frame.
+$\mathrm{obj}$ and $\mathrm{cls}$ are object confidence scores and classification scores.
+Our focus is to suppress the person’s object confidence score by optimizing the use of adversarial patches.
+IV-C Experimental Environment and Dataset
 For training and evaluating the proposed approach, we utilize the Inria Person dataset[75], which is specifically designed for pedestrian detection tasks.
@@
 By varying the angles ($0^{\circ}$, $30^{\circ}$, $60^{\circ}$, $90^{\circ}$) and distances (1 meter, 2 meters, 3 meters, 4 meters) between the human and the camera, we investigate the attack performance of MVPatch and other similar attack methods in a physical setting.
-IV-C Evaluation Metrics
+IV-D Evaluation Metrics
 To evaluate the naturalness and attack transferability of MVPatch, we employed the following evaluation metrics: mean Average Precision (mAP), Naturalness Score (NS), and Transferability Score (TS).
-mAP (mean Average Precision): The mAP is a widely used metric calculated by summing the average precision of models for each test sample and dividing it by the total number of test samples.
+mean Average Precision (mAP): The mAP is a widely used metric calculated by summing the average precision of models for each test sample and dividing it by the total number of test samples.
 It allows for comparisons of model performance across different datasets.
-The precision PREC is computed as the ratio of true positive samples divided by the sum of true positive and false positive samples.
-The recall REC is calculated as the ratio of true positive samples divided by the sum of true positive and false negative samples.
+The precision ‘PREC’ is computed as the ratio of true positive samples divided by the sum of true positive and false positive samples.
+The recall ‘REC’ is calculated as the ratio of true positive samples divided by the sum of true positive and false negative samples.
 Formally,
@@
 Through the equation mentioned above, we can obtain the mAP of the object detection models.
-NS (Naturalness Score): The NS is used to measure the similarity between the adversarial patch and the specified image, which is defined as
-$\mathrm{NS}=$ $\lambda\left(\frac{\mathrm{CSS}-\widehat{\mathrm{CSS}}_{\mathrm{NG}}}{\widehat{\mathrm{CSS}}_{\mathrm{S}}-\widehat{\mathrm{CSS}}_{\mathrm{NG}}}+\frac{\mathrm{CSS}-\widehat{\mathrm{CSS}}_{\mathrm{NR}}}{\widehat{\mathrm{CSS}}_{\mathrm{S}}-\widehat{\mathrm{CSS}}_{\mathrm{NR}}}\right)$ (12)
+Naturalness Score (NS): The NS is used to measure the similarity between the adversarial patch and the specified image, which is defined as
+$\mathrm{NS}=$ $\lambda\left(\frac{\mathrm{CS}-\widehat{\mathrm{CS}}_{\mathrm{NG}}}{\widehat{\mathrm{CS}}_{\mathrm{S}}-\widehat{\mathrm{CS}}_{\mathrm{NG}}}+\frac{\mathrm{CS}-\widehat{\mathrm{CS}}_{\mathrm{NR}}}{\widehat{\mathrm{CS}}_{\mathrm{S}}-\widehat{\mathrm{CS}}_{\mathrm{NR}}}\right)$ (26)
 $+\left(1-\lambda\right)\left(\frac{\mathrm{ED}-\widehat{\mathrm{ED}}_{\mathrm{NG}}}{\widehat{\mathrm{ED}}_{\mathrm{NG}}-\widehat{\mathrm{ED}}_{\mathrm{S}}}+\frac{\mathrm{ED}-\widehat{\mathrm{ED}}_{\mathrm{NR}}}{\widehat{\mathrm{ED}}_{\mathrm{NR}}-\widehat{\mathrm{ED}}_{\mathrm{S}}}\right),$
-where CSS represents the cosine similarity score between the specified image and adversarial patch and $\widehat{\text{CSS}}$ is the cosine similarity score between the specified image and adversarial patch of NG and NR images.
-ED denotes the Euclidean distance.
-NG and NR are gray-scale and random noise.
-S is an abbreviation for the source image.
-A higher naturalness score indicates that MVPatch has a stronger similarity to the specified image.
-TS (Transferability Score): The TS measures the transferable attack performance of the adversarial patch across diverse object detection models, which is defined as
+where CS denotes the cosine similarity score between the specified image S and the adversarial patch.
+$\widehat{\text{CS}}$ signifies the cosine similarity score between the specified image S and the adversarial patch of grayscale NG and random noise NR.
+Furthermore, ED represents the Euclidean distance.
+A higher naturalness score implies that the generated adversarial patch exhibits a stronger resemblance to the specified image.
+Transferability Score (TS): The TS measures the transferable attack performance of the adversarial patch across diverse object detection models, which is defined as
 $\mathrm{TS}=\frac{1}{N}\sum_{i=0}^{N}\left(\frac{\|D_{i}-\widehat{D}_{i}^{\mathrm{NG}}\|_{2}}{\widehat{D}_{i}^{\mathrm{NG}}}+\frac{\|D_{i}-\widehat{D}_{i}^{\mathrm{NR}}\|_{2}}{\widehat{D}_{i}^{\mathrm{NR}}}\right),$ (27)
-where $D_{i}$ denotes the mAP of object detection models detecting normal images while $\widehat{D}_{i}$ represents the mAP of object detection models detecting NG and NR images.
-NG and NR represent gray-scale and random noise.
-$||\cdot||_{2}$ is the L2 norm to constrain the distances between $D_{i}$ and $\widehat{D}_{i}$.
-The higher transferability score reveals that the adversarial patch has a higher transferable attack performance.
-We define
-$\mathrm{ASR}=\frac{1}{|\mathcal{D}_{\mathrm{test}}|}\sum_{\mathcal{I}\in\mathcal{D}_{\mathrm{test}}}\mathbbm{1}\{f(\mathcal{I^{*},\theta})\neq f(\mathcal{I},\theta)\},$ (14)
-as an indicator of the attack success rate, where $f(\cdot)$ represents the consequence of the models’ detection.
-$\mathcal{D}_{\mathrm{test}}$ denotes a set of benign images from test datasets.
-The ASR equals 1 when $f(\mathcal{I^{*},\theta})\neq f(\mathcal{I},\theta)$, and equals 0 otherwise.
+where $D_{i}$ denotes the mean Average Precision (mAP) of object detection models in identifying normal images, whereas $\widehat{D}i$ represents the mAP of object detection models in identifying NG (gray-scale) and NR (random noise) images.
+The symbols NG and NR respectively denote grayscale and random noise images.
+The term $||\cdot||_{2}$ refers to the L2 norm, which is employed to constrain the distances between $D_{i}$ and $\widehat{D}i$.
+A higher transferability score indicates that the adversarial patch demonstrates superior transferable attack performance.
+The Attack Success Rate (ASR) is defined as follows:
+$\mathrm{ASR}=\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{\mathcal{I}_{\mathrm{test}}}\mathbbm{1}\{f(\mathcal{I^{*},\theta})\neq f(\mathcal{I},\theta)\},$ (28)
+where $\mathcal{I}_{\mathrm{test}}$ denotes a set of benign images derived from test datasets, and $f(\cdot)$ represents the outcome of the models’ detection process.
+The ASR serves as an indicator of the attack success rate.
+It equals 1 if $f(\mathcal{I^{*},\theta})\neq f(\mathcal{I},\theta)$, and 0 otherwise.
 V Results
-This section presents empirical evidence that showcases the effectiveness of the proposed MVPatch through comprehensive evaluations conducted in both digital and physical domains.
-In the digital domain, we divided the experimental objects into meaningful adversarial patches and meaningless adversarial patches.
-In the physical domain, we assessed the patch attack performance from various angles and distances by utilizing a camera to capture images as input sources.
-The experimental results demonstrate that our method outperforms other similar adversarial patch attack methods, both in the digital and physical domains.
-In the last, we show the visual consequence of the experiment in the digital domain and physical domain.
-V-A Evaluation of Naturalness in the Digital World
-To assess the naturalness of the adversarial patch, we utilize the YOLO series models as threat models.
-The natural factor, represented by the parameter $\lambda$, is adjusted within the range of [1,3] to generate a variety of adversarial patches.
-The naturalness score (NS) is then calculated to quantify the naturalness of these patches.
-The generated adversarial patches with different naturalness scores are displayed in Fig.3.
-As we move from left to right on the coordinate axis, the naturalness score gradually increases, indicating a higher level of naturalness.
+TABLE I: Attack Performance and Comprehensive Evaluation with Diverse Objective Loss Functions
+Faster-RCNN SSD YOLOv2 YOLOv3 YOLOv3-tiny YOLOv4 YOLOv4-tiny YOLOv5 Mean mAP CS ED
+$\mathcal{L}_{\text{OBJ}_{1}}$ 42.07% 54.57% 30.26% 31.39% 13.49% 27.31% 22.74% 35.72% 32.20% 97.91% 0.0428
+$\mathcal{L}_{\text{OBJ}_{2}}$ 49.40% 61.08% 50.34% 37.47% 41.44% 52.64% 56.46% 56.46% 51.80% 97.91% 0.0419
+In this section, we present empirical evidence that underscores the efficacy of the proposed MVPatch through extensive evaluations conducted in both digital and physical domains.
+Initially, we illustrate the experimental results of various ensemble model strategies, demonstrating that our proposed ensemble model strategy exhibits superior performance compared to alternative approaches.
+Additionally, we evaluate naturalness by varying the natural factor $\lambda$.
+Within the digital domain, we categorize the experimental subjects into meaningful and meaningless adversarial patches.
+In the physical domain, we assess the performance of the patch attack from different angles and distances by employing a camera to capture images as input sources.
+The experimental outcomes reveal that our method surpasses other comparable adversarial patch attack methods in both digital and physical domains.
+Lastly, we present the visual outcomes of the experiments conducted in both the digital and physical domains.
+V-A Evaluation of Transferability with Diverse Ensemble Model Strategies
+Based on theoretical analysis discussed in Section III-C concerning the enhancement of transferability, the ensemble modeling technique can achieve superior adversarial transferability performance across multiple models when compared to a single model.
+There are, however, two primary approaches for leveraging the ensemble method: maximizing the outputs of neural networks and averaging the outputs of neural networks.
+We first investigate the utilization of two distinct loss functions for attacking object detection models, as outlined below:
 TABLE II: Experimental Consequence of Naturalness of MVPatch
@@
 $\lambda=3$ 39.69% 99.44% 1.99 81.01
-Explanation for Experimental Consequence: $\lambda=[1,3]$ represents the adversarial patches generated by diverse natural factors.
-SRC represents the source image.
-NoiseGrey and NoiseRandom are matrices with values of 0.5 and random values, respectively.
-We deploy the mAP (mean Average Precision), CS (Cosine Similarity), and ED (Euclidean Distance) to measure the attack performance and the similarity of the adversarial patch with the source image.
+$\mathcal{L}_{\mathrm{OBJ_{1}}}=\frac{1}{M}\sum_{i=1}^{M}f_{i}(\hat{x}),$ (29)
+$\mathcal{L}_{\mathrm{OBJ_{2}}}=\max\{f_{i}(\hat{x})\}^{M}_{i=1},$ (30)
+where $f(\cdot)$ denotes the object confidence score of the $i^{th}$ object detection model in the ensemble of object detection models (YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny).
+$\hat{x}\in\mathcal{I^{*}}$ is the image with adversarial patch.
+We compare the attack performance of diverse object confidence score loss functions on various object detection models.
+It is observed that the adversarial patch generated by $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ outperforms the one generated by $\mathcal{L}_{\mathrm{{OBJ}_{2}}}$ in terms of attacking the object detection model.
+The adversarial patches generated through diverse loss functions are then applied to eight object detection models (Faster-RCNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5 models) to assess their attack performance and similarity to specified images.
+The adversarial patch generated by $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ not only exhibits excellent attack capability (the average mAP of $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ is 32.20% and the average mAP of $\mathcal{L}_{\mathrm{{OBJ}_{2}}}$ is 51.80%) but also maintains a high level of naturalness.
+Consequently, we select $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ as our objective loss function.
 Figure 3: The naturalness score (NS) of adversarial patches with diverse $\lambda$ parameters. As the NS increases, the level of similarity between the source image and the generated image rises.
-Analysis and Conclusion for Experimental Consequence: As shown in Table II, we vary $\lambda$ from 1 to 3, resulting in an increase in the score of naturalness from 35.59 to 81.01.
-Simultaneously, the mAP increases from 32.30% to 39.69%.
-However, a higher mAP generally corresponds to greater accuracy of the object detection models and a lower attack performance of adversarial patches.
-There is a tradeoff between naturalness and attack performance, so it is not possible to obtain an adversarial patch with both higher naturalness and attack performance.
-We identify a threshold for the score of naturalness that helps us obtain the optimal adversarial patch with both naturalness and attack performance.
-Through an analysis of 20 experimental participants, it is generally agreed that when the score of naturalness is above 75 ($\lambda=2.5$), the generated adversarial patch would be considered more natural.
-Furthermore, if we continue to increase the natural factor $\lambda$, the gain in naturalness recognizable to the human eye will decrease.
-Therefore, we define $\lambda=2.5$ as the parameter of the loss function in the following experiments.
-The images with different scores of naturalness for the adversarial patches are shown in Fig.3.
+The results of the experiment are illustrated in Table I, where the Faster-RCNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5 models are included as threat models.
+The mean mAP represents the average value of the mean Average Precision (mAP) obtained from diverse models.
+CS corresponds to the Cosine Similarity and ED denotes the Euclidean Distance.
+V-B Evaluation of Naturalness with Diverse Natural Factors
+We assessed adversarial patches generated using YOLO models, adjusting the natural factor $\lambda$ within [1,3], and found that increasing $\lambda$ enhances naturalness, measured by the naturalness score (NS).
+The attack performance and similarity to the source image (SRC) were evaluated using mean Average Precision (mAP), Cosine Similarity (CS), and Euclidean Distance (ED).
+Additionally, NoiseGrey and NoiseRandom are matrices with values of 0.5 and random values, respectively.
+As shown in Table II, varying $\lambda$ from 1 to 3 increases the naturalness score (NS) from 35.59 to 81.01.
+Simultaneously, the mean Average Precision (mAP) rises from 32.30% to 39.69%.
+However, a higher mAP generally indicates greater accuracy of the object detection models, leading to a lower attack performance of the adversarial patches.
+This highlights a tradeoff between naturalness and attack performance, making it challenging to achieve both high naturalness and high attack performance in a single adversarial patch.
+Through analysis, we identify a threshold for naturalness that optimizes both aspects.
+Feedback from 20 experimental participants indicates that a naturalness score above 75 (achieved with $\lambda=2.5$) is generally perceived as more natural.
+Beyond this point, further increases in $\lambda$ result in diminishing gains in naturalness as perceived by the human eye.
+Therefore, we define $\lambda=2.5$ as the optimal parameter for the loss function in subsequent experiments.
+Figure 3 illustrates images of adversarial patches with different naturalness scores.
 TABLE III: Experimental Consequence of Transferability of Meaningful Adversarial Patches (mAP %)
-${}^{(MeaningfulAdversarialPatches)}$hreat Models Transfer Attack Models Black Box(TS $\uparrow$)
+(MeaningfulAdversarialPatches)Threat Models Transfer Attack Models Black Box(TS $\uparrow$)
 Faster RCNN SSD YOLOv2 YOLOv3 YOLOv3-tiny YOLOv4 YOLOv4-tiny YOLOv5
@@
 […]
-$\bm{{}^{\mathrm{(-)}}}$nsemble Model[1] 61.28 52.28 49.42 35.46 25.29 51.71 18.51 64.00 40
+(-)Ensemble Model[1] 61.28 52.28 49.42 35.46 25.29 51.71 18.51 64.00 40
 […]
@@
 Fig.6(b) illustrates the impact of varying distances between the camera and patches on the ASR of different attack methods, such as Natural Patch, AdvPatch, AdvTexture, Benign Image, and MVPatch.
-V-B Evaluation of Attack Transferability in the Digital domain
-To systematically investigate the attack performance of adversarial patches in the digital domain, we divide the experimental objects into two parts: meaningful adversarial patches and meaningless adversarial patches.
-Most research mainly focused on meaningless adversarial patches and few researchers investigated the meaningful and meaningless adversarial patches in the digital and physical domains.
-V-B1 Evaluation of Attack Transferability of Meaningful Adversarial Patches in the Digital domain
-We utilize the pre-trained models including YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny to generate diverse adversarial patches with meaningful content using the natural factor $\lambda=2.5$, as shown in Fig.4, 4, 4, 4, and 4.
-By combining these pre-trained models, we obtain ensemble attack adversarial patches, as shown in Fig.4 and 4.
-To compare the influence of person images and other images as adversarial patches for the object detection models, we use the sunflower image (Fig.4) as a control test for the person image (Fig.4).
-We input all the acquired adversarial patches into transfer attack models, including Faster R-CNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5, to calculate the transferability score and compare it with the Natural Patch[1].
-The comparable experimental results are shown in Table III.
-Explanation for Experimental Consequence: YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, YOLOv5, Faster R-CNN, and SSD are pre-trained object detection models.
-The ”Threat Models” column represents diverse adversarial patches generated by different models, including YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and the Ensemble Model, which combines YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny.
-The ”original image” refers to an image without an adversarial attack.”Grey-scale Noise” and ”Random Noise” separately indicate an image with padding of a value of 0.5 and an image with random values.
-The Black Box (TS) measures the transferability of adversarial attacks, with the transferability score (TS) indicating superior performance with a higher score.
-Evaluation and Analysis for Experimental Consequence: Firstly, we compare the transferability score of MVPatch (Ours) to that of Natural Patch[1] on individual object detection models.
-MVPatch performs better than Natural Patch on most of the models in the experiment.
-For the YOLOv4 model, the TS of MVPatch is approximately 33% higher than the TS of Natural Patch.
-Similarly, the excellent attack performance of MVPatch is observed on the other four object detection models (YOLOv2, YOLOv3, YOLOv3-tiny, and YOLOv4-tiny).
-On the other hand, the attack performance of MVPatch on the ensemble model is surpassed by Natural Patch, with the TS of MVPatch on the ensemble model being approximately 16% higher than the TS of Natural Patch.
-From the results of the meaningful adversarial patch experiment, it can be seen that the proposed algorithm for meaningful adversarial patches, MVPatch, achieves superior performance compared to the similar algorithm, Natural Patch, in terms of attack transferability.
-V-B2 Evaluation of Attack Transferability of Meaningless Adversarial Patches in the Digital domain
-We adjust the natural factor of $\lambda=0$ and utilize the joint pre-trained models, including YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny, to generate meaningless adversarial patch MVPatch.
-In comparison, we consider AdvPatch[2], AdvTexture[3], and AdvCloak[32] as subjects for the experiment on meaningless adversarial patches.
-We use all the obtained adversarial patches as input for Faster R-CNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5, calculate the transferability scores, and compare them with those of AdvPatch [2], AdvTexture[3], and AdvCloak[32].
-The comparable experimental results are presented in TableIV.
-Explanation for Experimental Consequence: YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, YOLOv5, Faster R-CNN, and SSD are pre-trained object detection models.
-The ”Meaningless Adversarial Patches” column represents various meaningless adversarial patches generated by diverse attack algorithms.
-The Black Box (TS) measures the transferability of adversarial attacks, with a higher transferability score indicating better performance.
-TS stands for the transferability score.
-Evaluation and Analysis for Experimental Consequence: As illustrated in Table IV, the transferability score (TS) of MVPatch is approximately 40% higher than that of AdvCloak.
+V-C Evaluation in the Digital Domain
+To systematically investigate the attack performance of adversarial patches, we categorize the experiments into two parts: meaningful and meaningless adversarial patches.
+Most research has primarily focused on meaningless patches, while the exploration of both meaningful and meaningless patches in digital and physical domains remains limited.
+V-C1 Meaningful Adversarial Patches
+We utilize pre-trained models, including YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny, to generate diverse adversarial patches with meaningful content using a natural factor of $\lambda=2.5$, as shown in Figs.
+4 to 4.
+By combining these models, we create ensemble attack adversarial patches (Figs.
+4 and 4).
+To compare the effects of person images and other images, we use a sunflower image as a control for the person image.
+All generated patches are tested against transfer attack models, including Faster R-CNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5, to calculate transferability scores (TS) and compare them with Natural Patch [1].
+The results are presented in Table III.
+The ”Threat Models” column indicates adversarial patches generated by various models, while ”Grey-scale Noise” and ”Random Noise” refer to images padded with a value of 0.5 and random values, respectively.
+Higher TS indicates better transferability of the adversarial attacks.
+Analysis and Discussion: We compare the transferability score (TS) of MVPatch (Ours) to that of Natural Patch [1] on individual object detection models.
+MVPatch performs better than Natural Patch on most models in the experiment.
+For the YOLOv4 model, the TS of MVPatch is approximately 33% higher than that of Natural Patch.
+Similarly, MVPatch demonstrates excellent attack performance on other object detection models, including YOLOv2, YOLOv3, YOLOv3-tiny, and YOLOv4-tiny.
+Additionally, MVPatch surpasses Natural Patch in terms of attack performance on the ensemble model, with the TS of MVPatch being approximately 16% higher than that of Natural Patch.
+The results of the meaningful adversarial patch experiment clearly indicate that the proposed algorithm, MVPatch, achieves superior performance compared to the similar algorithm, Natural Patch, in terms of attack transferability.
+V-C2 Meaningless Adversarial Patches
+We set the natural factor to $\lambda=0$ and employ joint pre-trained models, including YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny, to generate the meaningless adversarial patch.
+We compare this with AdvPatch [2], AdvTexture [3], and AdvCloak [33].
+All generated patches are tested on Faster R-CNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5 to calculate transferability scores (TS), which are then compared with the benchmarks.
+The results are summarized in Table IV.
+The ”Meaningless Adversarial Patches” column lists patches generated by various attack algorithms, while the Black Box (TS) indicates transferability, with higher scores reflecting superior performance.
+Analysis and Discussion: As illustrated in Table IV, the transferability score (TS) of MVPatch is approximately 40% higher than that of AdvCloak.
 Compared to AdvPatch and AdvTexture, the TS of MVPatch is approximately 20% higher.
@@
 To sum up, in terms of meaningless adversarial patches, the proposed algorithm, MVPatch, exhibits higher attack transferability compared to similar algorithms.
-V-C Evaluation of Attack Performance of Adversarial Patches in the Physical domain
+V-D Evaluation in the Physical Domain
 To systematically investigate the attack performance of adversarial patches in the physical domain, we divide the experimental conditions into two parts: diverse angles and diverse distances.
@@
 The images at the bottom are adversarial patches and mosaics are added after detection, to avoid privacy disclosure for volunteers.
-V-C1 Evaluation of Attack Performance of Diverse Adversarial Patches in the Physical Domain
-The volunteers hold various adversarial patches, including AdvPatch, AdvTexture, a benign image, Natural Patch, and the proposed MVPatch, against their chests.
-The patches are positioned at an angle of {$0^{\circ},30^{\circ},60^{\circ},90^{\circ}$} and a distance of {1, 2, 3, 4} meters.
-We use an iPhone 11 to capture the images with the adversarial patches from diverse distances and angles in the physical world.
-In total, we collected 9681 images.
-To preserve the pixel details of the adversarial patches, we view them using an iPad Pro.
-The attack success rates and some evaluation metrics of different adversarial patches are shown in Table V, and sample patches are displayed in Fig.5.
-Explanation for Experimental Consequence:The AdvPatch[2], and AdvTetxure[3] are both meaningless adversarial patches and the Natural Patches[1] and proposed MVPatches are meaningful patches.
-The Benign Image refers to an image without adversarial examples.
-ASR is short for attack success rate.
-Number of Images represents the number of experimental images to evaluate the attack performance.
-Stealthiness Attack means that the attack has a more naturalness score.
-Physical Attack denotes that the attack can be applied in the physical world.
-Transferable Attack represents that the attack has transferable attack ability.
-A lightweight model means that the model, that generates the adversarial patches, has low computational cost.
-Evaluation and Analysis for Experimental Consequence: As Table V illustrates, both AdvPatch(Fig.5) and AdvTetxure(Fig.5) do not have much more attack performance on the YOLOv2 model in the physical world and the ASRs are 13.57% and 12.72%, respectively.
+V-D1 Comparision with Other Adversarial Patches
+Table V presents the empirical results of various adversarial patches within the physical domain, including AdvPatch, AdvTexture, a benign image, Natural Patch, and the proposed MVPatch.
+Volunteers held these patches against their chests at angles of $0^{\circ},30^{\circ},60^{\circ},90^{\circ}$ and distances of 1, 2, 3, 4 meters.
+Using an iPhone 11, we captured a total of 9,681 images from various angles and distances, and reviewed them on an iPad Pro to ensure pixel detail preservation.
+The table evaluates key metrics such as Attack Success Rate (ASR) on the YOLOv2 model, number of images, stealthiness, physical applicability, transferability, and computational cost.
+AdvPatch and AdvTexture are classified as meaningless adversarial patches, while Natural Patch and MVPatch are meaningful.
+The benign image contains no adversarial features.
+ASR measures the attack success rate, ”Number of Images” indicates the experimental samples, ”Stealthiness” reflects naturalness, ”Physical Attack” confirms real-world applicability, ”Transferable Attack” indicates cross-domain effectiveness, and ”Lightweight Model” denotes low computational cost.
+The proposed MVPatch demonstrates superior performance across all evaluated metrics.
+Analysis and Discussion: As Table V illustrates, both AdvPatch(Fig.5) and AdvTetxure(Fig.5) do not have much more attack performance on the YOLOv2 model in the physical world and the ASRs are 13.57% and 12.72%, respectively.
 The Natural Patch(Fig.5) has higher ASR and its value of ASR is 19.43%, but the other Natural Patch(Fig.5) does not perform as well as Fig.5.
@@
 Consequently, we ensure the MVPatch has more attack performance than others in the physical world.
-V-C2 Evaluation of Attack Performance from Diverse Angles in the Physical Domain
-We collected 6846 images from various angles ($0^{\circ}$, $30^{\circ}$, $60^{\circ}$, $90^{\circ}$) to assess the impact of different angles on the attack performance of adversarial patches, including the Natural Patch, AdvPatch, AdvTexture, Benign Image, and our MVPatch.
-To keep the fairness of the experiment, we change angles from {$0^{\circ}$, $30^{\circ}$, $60^{\circ}$, $90^{\circ}$} with the same distance (2 meters).
-The volunteer rotated while the adversarial patch rotated in the opposite direction, ensuring that the patch always faced the camera.
-We refer to this setup as ”with Rotation”.
-The experimental results are presented in Fig.6(a).
-Explanation for Experimental Consequence: The light blue dotted line with squares represents the MVPatch with Rotation, while the red dotted line with circles represents the Benign Image with Rotation.
-The dark blue dotted line with triangles represents the AdvPatch with Rotation, while the green dotted line with triangles represents the AdvTexture with Rotation.
-The black dotted line with squares represents the Natural Patch with Rotation.
+V-D2 Diverse Angles
+We collected 6,846 images from various angles ($0^{\circ}$, $30^{\circ}$, $60^{\circ}$, $90^{\circ}$) to assess the impact of angle variation on the attack performance of different adversarial patches, including Natural Patch, AdvPatch, AdvTexture, Benign Image, and our MVPatch.
+To ensure experimental fairness, angles were adjusted to $0^{\circ}$, $30^{\circ}$, $60^{\circ}$, $90^{\circ}$ while maintaining a constant distance of 2 meters.
+The volunteer rotated, with the adversarial patch rotating in the opposite direction to ensure it always faced the camera, a setup referred to as ”with Rotation.” The experimental results are presented in Fig.6(a).
+The light blue dotted line with squares represents the MVPatch with Rotation; the red dotted line with circles represents the Benign Image with Rotation; the dark blue dotted line with triangles represents the AdvPatch with Rotation; the green dotted line with triangles represents the AdvTexture with Rotation; and the black dotted line with squares represents the Natural Patch with Rotation.
 The X-axis represents the angles, while the Y-axis represents the attack success rate.
-Evaluation and Analysis for Experimental Consequence: The effectiveness of adversarial patches with rotation diminishes as the angle of rotation increases.
+Analysis and Discussion: The effectiveness of adversarial patches with rotation diminishes as the angle of rotation increases.
 At an angle of $0^{\circ}$, the ASR of the MVPatch can reach 39.5%, and other patches perform well at the angle of $0^{\circ}$ except for the AdvTexture.
@@
 Fig.6(a) also demonstrates that the MVPatch outperforms other attack methods in terms of angle variations (at the angles of $0^{\circ}$, $30^{\circ}$, and $60^{\circ}$).
-V-C3 Evaluation of Attack Performance from Diverse Distance in the Physical Domain
-A total of 3984 images are collected at various distances (1 meter, 2 meters, 3 meters, and 4 meters) to assess how different distances impact the attack performance of adversarial patches, specifically MVPatch, Natural Patch, AdvPatch, AdvTexture, and Benign Image.
-To keep the fairness of the experiment, we change distances from {1, 2, 3, 4} meters with the same angle ($0^{\circ}$).
-Fig.6(b) presents the results.
-Explanation for Experimental Consequence: MVPatch is represented by the light blue dotted line with squares.
-Natural Patch is represented by the black dotted line with squares.
-AdvPatch is represented by the dark blue dotted line with triangles.
-AdvTexture is represented by the green dotted line with triangles.
-The benign Image is represented by the red dotted line with circles.
-The X-axis represents the attack success rate, and the Y-axis represents the distance from the volunteer with the adversarial patch to the camera.
-Evaluation and Analysis for Experimental Consequence: As Fig.6(b) shows, the adversarial patches generated by MVPatch and AdvPatch exhibit high levels of aggression when the volunteer is at a distance of one meter from the camera.
+V-D3 Diverse Distance
+A total of 3,984 images were collected at various distances (1 meter, 2 meters, 3 meters, and 4 meters) to evaluate the impact of distance on the attack performance of adversarial patches, including MVPatch, Natural Patch, AdvPatch, AdvTexture, and Benign Image.
+To ensure experimental fairness, distances were varied 1, 2, 3, 4 meters while maintaining a constant angle of $0^{\circ}$.
+The results are presented in Fig.6(b).
+The light blue dotted line with squares represents the MVPatch, the black dotted line with squares represents the Natural Patch, the dark blue dotted line with triangles represents the AdvPatch, the green dotted line with triangles represents the AdvTexture, and the red dotted line with circles represents the Benign Image.
+The X-axis denotes the attack success rate, and the Y-axis indicates the distance from the volunteer with the adversarial patch to the camera.
+Analysis and Discussion: As Fig.6(b) shows, the adversarial patches generated by MVPatch and AdvPatch exhibit high levels of aggression when the volunteer is at a distance of one meter from the camera.
 The ASR of MVPatch is 71.67% and the ASR of AdvPatch is 63.46% at a distance of 1 meter.
@@
 Similarly, the attack weakens as the volunteer moves closer.
-V-D Illustraton of Experimental Consequence
-This section provides experimental results in both the digital and physical domains to demonstrate our proposed approach.
+V-E Visualization
+This section provides experimental results in both the digital and physical domains to visualize our proposed approach.
 The green rectangle represents a successful detection of a person by the object detector, while the red rectangle indicates a failure to detect the person class.
@@
 Fig.7 displays the results of digital experiments, while Fig.8 depicts the results of physical experiments.
-VI Conclusion
-In this paper, we propose a dual-perception-based attack framework.
-This framework comprises a model-perception degradation method and a human-perception improvement method.
-The aim is to create vivid and aggressive adversarial patches, which we refer to as MVPatch, for person detection models suitable for real-world implementation.
-In terms of the model-perception degradation method, we exploit the ensemble attack on various object detectors to enhance the transferability of the adversarial patch.
-In terms of the human-perception improvement method, we introduce a lightweight approach for visual similarity measurement designed to make the adversarial patch less noticeable.
-Through extensive qualitative and quantitative experiments, we compare different approaches and find that minimizing the object loss yields the most effective patches while maintaining competitive attack performance compared to similar methods.
-We evaluate the transferability and naturalness of the crafted adversarial patches in both digital and physical domains.
-Extensive experiments demonstrate that the proposed MVPatch performs remarkable stealthiness and transferability with few computational costs.
-However, there are still limitations in terms of factors such as angles and distance that can affect the performance of adversarial attacks.
-In future work, we aim to explore approaches to designing adversarial patches that can minimize the influence of angle and distance and maximize attack performance in both digital and physical domains.
+VI Conclusion and Limitations
+In this paper, we propose a Dual-Perception-Based Framework (DPBF) that includes a Model-Perception-Based Module (MPBM), a Human-Perception-Based Module (HPBM), and regularization terms.
+Our objective is to create vivid and aggressive adversarial patches, termed More Vivid Patch (MVPatch), designed to camouflage individuals in detection models for real-world applicability.
+The MPBM leverages ensemble attacks across various object detectors to enhance the transferability of adversarial patches, demonstrating improved generalization and stability through generalization error analysis.
+Meanwhile, the HPBM introduces a lightweight approach for visual similarity measurement, which makes the adversarial patches less noticeable and incorporates additional transformations to further enhance transferability while preserving stealthiness.
+Regularization terms are employed to increase the practicality of the generated patches in physical environments.
+We also introduce metrics for naturalness and transferability to provide an unbiased assessment of the adversarial patches.
+Extensive qualitative and quantitative experiments reveal that minimizing our proposed objective function results in the most effective patches, maintaining competitive performance compared to existing methods.
+The MVPatch demonstrates exceptional stealthiness and transferability across both digital and physical domains, achieving these results with minimal computational cost.
+Limitations.
+Despite the advancements presented, our approach has several limitations.
+Factors such as angles and distances can still affect the effectiveness of adversarial attacks.
+Additionally, the vulnerability of adversarial patches against state-of-the-art object detection models remains inadequately explored.
+Future work will address these issues by developing methods to minimize the impact of angle and distance, thereby optimizing attack performance in both digital and physical contexts.
+We also intend to apply our technique to a wider range of off-the-shelf object detection models to evaluate their susceptibility to adversarial patches.
+Furthermore, we plan to explore more robust defense mechanisms to mitigate the risk of adversarial attacks in both digital and physical environments.